### 30 Mar 2026 v1

# On the Complexity of Determinations

Abstract
Classical complexity theory measures the cost of computing a function, but many computational tasks require committing to one valid output among several. We introduce determination depth—the minimum number of sequential layers of irrevocable commitments needed to select a single valid output—and show it is an orthogonal axis measuring the cost of committing. We exhibit relational tasks whose commitments are constant-time table lookups yet require exponential parallel width to compensate for any reduction in depth. A conservation law couples determination and computational depth: enriching the commitment basis trades layers for step complexity, but the total sequential cost is bounded below. For circuit-encoded specifications, the resulting depth hierarchy captures the polynomial hierarchy ($\Sigma_{2k}^{P}$-complete for each fixed $k$, PSPACE-complete for unbounded $k$). Determination depth and computational depth are orthogonal: neither subsumes the other. Stable matching witnesses the independence sharply—finding a stable matching is in P, yet every finite determination depth arises as the rotation-poset height of some instance. In the distributed setting, the framework recovers the Halpern–Moses common-knowledge impossibility.
1 Introduction
Classical complexity theory measures the cost of computing a function—producing the unique correct output. But many tasks require resolving a relation: choosing one admissible outcome from a set, as when a text generator selects one coherent continuation among many, or a distributed system must agree on one valid output among several. We show that this distinction creates a new, orthogonal axis of complexity. Classical hard functions are computationally expensive but have no cost on this new axis. However, there exist relational tasks in which every step is a constant-time table lookup, yet any strategy using fewer than $k$ sequential layers requires exponential width (exponentially many parallel candidates), even with unbounded parallelism within each candidate. The bottleneck is not computation, nor communication, but irrevocable commitment. This cost is invisible to classical complexity measures—circuit depth, communication complexity, adaptive queries—because they track the cost of computing or distributing an answer, not the cost of committing to one.
We call this choice process a determination: a collection of irrevocable commitments that narrow the admissible set until a single outcome remains. Commitments whose order affects which outcomes survive must be sequenced into layers; those whose order does not matter can be applied in parallel. The determination depth of a specification is the minimum number of layers needed to resolve it—the intrinsic cost of committing.
We establish three main results. (1) Exponential separation (Section 3): the depth–width tradeoff is exponential, generalizing pointer chasing [NW93, MYZ25] from functions to relations. (2) Conservation law (Section 4): enriching the commitment basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below—choices can be bundled into richer steps, but the burden does not disappear. The bound is tight. Under a commutative basis in the online setting, a supporting oracle characterization shows that determination depth equals the minimum number of observations of the growing history, regardless of computational power between observations. (3) Polynomial-hierarchy characterization (Section 5): for circuit-encoded specifications, “is depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete for each fixed $k$ and PSPACE-complete for unbounded $k$, certifying determination depth as a semantic measure well-calibrated against the classical hierarchy.
The exponential separation construction is autoregressive: each commitment is a single irrevocable choice, and the depth-$k$ lower bound says that $k$ sequential choices are unavoidable regardless of parallelism. This connects determination depth to chain-of-thought reasoning, where each token of sequential inference is a commitment layer (Appendix D.3). Stable matching provides a complementary structural connection: every finite determination depth arises as the rotation-poset height of some instance (Appendix D.1).
The appendices apply the framework to several classical settings, decomposing known costs into intrinsic determination depth and circumstantial model assumptions, and in several cases revealing new structure that existing models do not capture. Each determination layer carries a concrete physical cost: a coordination round in distributed systems, a token of inference in language models, a round of play in a game. Existing models can hide this cost by assuming away commitments, or inflate it by charging for operations that resolve no ambiguity.
The appendices exhibit both phenomena. BSP round complexity overcharges by counting communication rounds that involve no commitment, and undercharges by assuming fixed membership. In extensive-form games, determination depth measures the game’s strategic depth—the number of moves where a player must break a tie among equally-optimal options—which can be strictly less than the game-tree depth; one player can influence the other’s strategic depth, creating a new dimension of strategic interaction (Appendix D.2). In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge [HM90], with the conservation law and exponential separation providing quantitative predictions beyond the binary threshold.
2 Framework
We model both online and offline settings via histories—partially ordered sets of events that represent everything that has occurred up to a given point.
Definition 1 (History).
A history is a partially ordered set $H=(E,\to)$ of events, where $\to$ is interpreted as causal precedence [LAM78]. Events are of two kinds: environment events (inputs, messages, computation steps), which are given, and commitment events (defined below), which are chosen.
The results below are parametric in a class $\mathcal{H}$ of histories under consideration.
Definition 2 (History Extension).
For histories $H_{1}=(E_{1},\to_{1})$ and $H_{2}=(E_{2},\to_{2})$ in $\mathcal{H}$, we write $H_{1}\sqsubseteq H_{2}$ if $E_{1}\subseteq E_{2}$ and $H_{1}$ is a downward-closed sub-poset of $H_{2}$ (i.e., whenever $e\in E_{1}$ and $e^{\prime}\to e$ in $H_{2}$, then $e^{\prime}\in E_{1}$).
[…]
Definition 3 (Specification).
A specification is a mapping $\mathsf{Spec}:\mathcal{H}\to 2^{O}$ assigning to each history a set of admissible outcomes. We call $\mathsf{Spec}(H)$ the admissible set at $H$.
A specification may admit multiple outcomes for the same history; resolving this ambiguity is the cost that determination complexity measures.
Definition 4 (Commitment).
A commitment $\varphi$ is a constraint on outcomes. We write $H\cdot\varphi$ for the history obtained by adding a commitment event $\varphi$ after all maximal events of $H$ (i.e., $e\to\varphi$ for every $e$ with no successor in $H$), giving $H\sqsubseteq H\cdot\varphi$. Adding a commitment can only narrow the admissible set (shrinkage): $\mathsf{Spec}(H\cdot\varphi)\subseteq\mathsf{Spec}(H)$.
In the distributed setting, each agent’s commitments succeed only its own local maximal events, not the global set; Appendix B develops this extension fully.
Each commitment event $\varphi$ in a history carries its context: the prefix of the history up to $\varphi$ determines the admissible set at which $\varphi$ takes effect. Along any chain of extensions, environment events may enlarge or shrink the admissible set, but each commitment event can only shrink it (shrinkage condition). A relational specification cannot be resolved to a function by environment events alone; commitments are required, and the cost of those commitments is what determination depth measures (Section 2.1).
Definition 5 (Determination).
The determination $D(H)$ of a history $H$ over a commitment basis $\Phi$ is the subsequence $\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ of commitment events in $H$, listed in causal order. Its cost is $\mathsf{cost}(D(H))\triangleq m$.
The commitment events in any single-agent history are totally ordered (each follows all maximal events of its prefix), so the determination is a well-defined sequence. Writing $H_{j}$ for the prefix of $H$ through $\varphi_{j}$, shrinkage accumulates: $\mathsf{Spec}(H_{j})\subseteq\mathsf{Spec}(H_{j-1})$ for each $j$. A determination is valid if each $\varphi_{j}$ preserves nonemptiness: $\mathsf{Spec}(H^{\prime})\neq\emptyset$ for every $H^{\prime}\sqsupseteq H_{j}$. A specification is resolved if $|\mathsf{Spec}(H)|=1$ for all $H$; a history resolves $\mathsf{Spec}$ if its determination leaves $|\mathsf{Spec}(H)|=1$.
In the offline setting, environment events precede all commitment events, so the determination is a single consecutive block. In the online setting, environment events may interleave with commitments, breaking the determination into separate runs; the effect of a commitment depends on which environment events have occurred, so the online setting requires an adaptive strategy (formalized in Section A).
Three forces on a history.
A history evolves under three distinct forces: the environment extends the history with new events (input arrivals, message deliveries); commitments extend the history with events that irrevocably narrow the admissible set; and a determination strategy selects which commitment events to append, based on the history observed so far. A determination strategy is a rule $\sigma:\mathcal{H}\to\Phi^{*}$ mapping each observable history to a sequence of commitments from the basis. It is a semantic object: it specifies which commitments are made and when in history, but says nothing about how they are computed. A protocol, algorithm, or agent is a realization of a determination strategy; the complexity results in this paper are stated at the strategy level.
Commutation and depth.
Commitments $\varphi$ and $\psi$ commute with respect to $\mathsf{Spec}$ at $H$ if the order of extension does not matter: $\mathsf{Spec}(H\cdot\varphi\cdot\psi)=\mathsf{Spec}(H\cdot\psi\cdot\varphi)$. Commutation is relative to a specification and may change as the history grows: $\varphi$ and $\psi$ may commute at $H$ but not at an extension $H^{\prime}\sqsupseteq H$, or vice versa. This history-dependence does not arise in classical trace theory, where independence is static, and it is what makes non-commuting commitments impose irreducible sequencing. (A history records one linearization of the commitments; commutation ensures the choice of linearization is immaterial.) The depth of a determination is the total number of commuting layers across all runs of consecutive commitments, where commitments within each layer commute pairwise with respect to the specification at that point in the history. The full algebraic development appears in Section A.
Definition 6 (Determination cost and determination depth).
A determination strategy $\sigma$ resolves $\mathsf{Spec}$ if, for every history $H$ arising under $\sigma$, $|\mathsf{Spec}(H)|=1$ once all of $\sigma$’s commitments have been applied. Define $\mathsf{Cost}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;\mathsf{cost}(D(H))$ and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;0pt(D(H))$, where $\sigma$ ranges over strategies that resolve $\mathsf{Spec}$ and $H$ ranges over histories arising under $\sigma$; or $\infty$ if no resolving strategy exists. In the offline setting (a single history with no environment extensions), the max is trivial and the definition reduces to a min over determinations.
$\mathsf{Cost}_{\Phi}(\mathsf{Spec})$ is the determination cost and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ the determination depth. Both are properties of the specification, not of any particular strategy—the strategy is the witness, just as a circuit witnesses the depth of a function—and they are the semantic analogues of circuit size and circuit depth: cost counts commitments, depth counts sequential layers. In the online setting, the max over $H$ gives the environment adversarial control bounded by the history class $\mathcal{H}$; this is a worst-case measure, and the results in this paper focus on this setting.
2.1 Commitment bases
Determination depth depends on the commitment basis $\Phi$. We begin with the canonical minimal basis and then identify three properties it satisfies that generalize to richer settings.
Definition 7 (Atomic basis).
For a history $H$ and outcome $o$, the atomic commitment excluding $o$ at $H$ is the event $\varphi$ that, when appended to $H$, excludes exactly $o$: $\mathsf{Spec}(H\cdot\varphi)=\mathsf{Spec}(H)\setminus\{o\}$, and $o\notin\mathsf{Spec}(H^{\prime})$ for every $H^{\prime}\sqsupseteq H\cdot\varphi$. It has no effect when appended to a history incomparable to $H$. The atomic basis contains one such commitment per pair $(H,o)$.
Atomic commitments are the finest-grained irrevocable commitments: each permanently excludes a single outcome at a single history. The permanence condition ($o$ never reappears at any extension) is what makes the exclusion irrevocable; it is satisfied by all specifications studied in this paper.
In online settings, depth under the atomic basis arises from forward validity constraints: two atomic commitments made at history $H$ cannot be applied in the same layer if their joint effect empties the admissible set at an extension of $H$, as in the following example:
Example 1 (Three-valued consensus).
Let $O=\{a,b,c\}$ with $\mathsf{Spec}(H_{0})=\{a,b,c\}$, and suppose the environment may extend $H_{0}$ to a history $E\sqsupseteq H_{0}$ with $\mathsf{Spec}(E)=\{a,b\}$. The refined admissible sets $\mathsf{Spec}(H_{0}\cdot\varphi_{\neg a})$ and $\mathsf{Spec}(E\cdot\varphi_{\neg a})$ are both non-empty ($\{b,c\}$ and $\{b\}$ respectively), so $\varphi_{\neg a}$ is valid at $H_{0}$. But the joint refinement $\mathsf{Spec}(E\cdot\varphi_{\neg a}\cdot\varphi_{\neg b})$ is empty: $\{a,b\}\setminus\{a,b\}=\emptyset$. Symmetrically, if the environment may also extend $H_{0}$ to $E^{\prime}$ with $\mathsf{Spec}(E^{\prime})=\{b,c\}$ and to $E^{\prime\prime}$ with $\mathsf{Spec}(E^{\prime\prime})=\{a,c\}$, then no pair of atomic commitments at $H_{0}$ is jointly valid, so resolution requires two layers: commit, observe the environment, then commit again.
Definition 8 (Intrinsic determination depth).
The intrinsic determination depth of a specification $\mathsf{Spec}$ is $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ where $\Phi$ is the atomic basis.
Intrinsic depth measures the irreducible layering forced by forward validity constraints alone—the “bare” sequential cost before any computational or architectural assumptions are introduced. In the offline setting, forward validity constraints vanish and intrinsic depth collapses; the conservation law (Theorem 3) shows that the sequential cost reappears as computation within layers.
Properties of the atomic basis.
The atomic basis satisfies three properties that we now name for use in the general theory.
Definition 9 (Pointwise basis).
A commitment $\varphi$ is pointwise if it acts as an outcome-by-outcome filter: whether $o\in\mathsf{Spec}(H\cdot\varphi)$ depends only on $o$ and $H$, not on the rest of $\mathsf{Spec}(H)$. A pointwise basis consists entirely of pointwise commitments. A commitment whose effect on one outcome depends on the presence of another is non-pointwise.
Pointwise commitments always commute (each outcome is filtered independently, so application order is irrelevant). Atomic commitments are pointwise, but the pointwise class is strictly larger: a commitment that excludes multiple outcomes simultaneously (e.g., “keep only tuples with $v_{1}=3$”) is pointwise but not atomic. Commutativity can also hold for non-pointwise commitments:
Definition 10 (Commutative basis).
A commitment basis $\Phi$ is commutative if every pair of commitments in $\Phi$ commutes at every history for every specification: $\mathsf{Spec}(H\cdot\varphi\cdot\psi)=\mathsf{Spec}(H\cdot\psi\cdot\varphi)$ for all $\varphi,\psi\in\Phi$, all $\mathsf{Spec}$, and all $H\in\mathcal{H}$.
Definition 11 (Constant-depth basis).
A commitment basis $\Phi$ is constant-depth for $\mathsf{Spec}$ if each commitment in $\Phi$ is computable in $O(1)$ circuit depth from a standard polynomial-size encoding of the current history.
The atomic basis is both commutative and constant-depth, making it the canonical minimal basis for the online setting where commitments are made before all information has arrived. Under a constant-depth basis, every unit of determination depth corresponds to one unit of sequential time; under a richer basis, some of that time is absorbed into the computation needed to evaluate each commitment. The pointwise, commutative, and constant-depth properties each appear in multiple results: pointwise bases arise in the exponential separation (Section 3), the PH characterization (Section 5), and the NP-hardness of computing depth; commutative bases in the oracle characterization (Section 4.2) and the distributed extension (Appendix B); and constant-depth bases in the conservation law (Section 4.1).
3 Exponential Depth–Width Separation
We exhibit a family of constrained generation tasks—modeling autoregressive text generation—in which every commitment is a constant-time table lookup, yet any strategy using $d^{\prime}<k$ layers requires width at least $(m/s)^{k-d^{\prime}}$.
3.1 The constrained generation task
Fix positive integers $k$ (number of positions), $m$ (domain size), and $s$ (constraint sparsity), with $1\leq s\leq m$. The task has k1k-1 links $1\to 2\to\cdots\to k$, each constraining a position’s value given its predecessor’s.
Definition 12 ($k$-position constraint chain).
A $k$-position constraint chain over domain $[m]$ with sparsity $s$ is a sequence of constraint functions $P_{1}\subseteq[m]$ and $P_{\ell}\subseteq[m]\times[m]$ for $\ell=2,\ldots,k$, such that $|P_{1}|=s$ and, for each $\ell\geq 2$ and each $a\in[m]$, the successor set $P_{\ell}(a,\cdot)\triangleq\{b\in[m]:(a,b)\in P_{\ell}\}$ satisfies $|P_{\ell}(a,\cdot)|=s$.
Definition 13 ($k$-position generation task).
Given a $k$-position constraint chain $(P_{1},\ldots,P_{k})$ (encoded as environment events in an offline history), the constrained generation task is to output any admissible tuple in
$\mathsf{Spec}_{\mathsf{cg}}(P_{1},\ldots,P_{k})\;\triangleq\;\bigl\{\,(v_{1},\ldots,v_{k})\in[m]^{k}\;\bigm|\;v_{1}\in P_{1}\text{ and }(v_{\ell-1},v_{\ell})\in P_{\ell}\text{ for }\ell=2,\ldots,k\,\bigr\}.$
The specification is relational ($s^{k}$ admissible tuples when $s\geq 2$) and models autoregressive generation: each position is a token, $[m]$ is the vocabulary, and $P_{\ell}$ encodes local coherence constraints.
Commitment basis.
For each position $\ell\in\{1,\ldots,k\}$ and value $v\in[m]$, the commitment $\varphi_{\ell,v}$ filters to tuples with $v_{\ell}=v$. These are pointwise, but the links introduce dependency: committing to $v_{\ell}$ before $v_{\ell-1}$ may violate a constraint. Each commitment is a constant-time table lookup.
Observation 1 (Sequential strategy).
Committing to one position per layer in order $v_{1},v_{2},\ldots,v_{k}$ always produces a valid tuple (each $v_{\ell}\in P_{\ell}(v_{\ell-1},\cdot)$ exists by sparsity), using $k$ layers and width $1$.
3.2 Strategy Model and Lower Bound
We formalize what a strategy with fewer than $k$ layers can do.
Definition 14 ($d^{\prime}$-layer strategy with width $w$).
A $d^{\prime}$-layer strategy with width $w$ partitions the $k$ positions into $d^{\prime}$ groups $S_{1},\ldots,S_{d^{\prime}}$ and produces a single history carrying $w$ parallel candidate tuples, each contributing $k$ commitments (one per position) distributed across the $d^{\prime}$ layers. In layer $r$, for each candidate, the strategy commits to values at all positions in $S_{r}$; these values may depend on the full input and on earlier-layer commitments, but not on same-layer or later-layer commitments. The strategy selects one of the $w$ candidates as its output; it resolves the specification if the output lies in $\mathsf{Spec}_{\mathsf{cg}}$.
The model is strictly more permissive than autoregressive generation (which fixes left-to-right order and width $1$): it allows arbitrary layer assignments, arbitrary within-layer parallelism, and $w$ independent candidates. The only constraint is that commitments within a layer cannot depend on same-layer outcomes—the defining property of parallel execution—so any lower bound here applies to every parallel strategy that respects this constraint, including beam search, speculative decoding, and diffusion-style refinement.
Each group $S_{r}$ is a commuting layer in which the $w$ candidates each commit to $|S_{r}|$ positions in parallel. When $d^{\prime}<k$, some adjacent positions share a layer and cannot observe each other’s values; the lower bound counts these blind spots (uninformed links) and shows that $w$ grows exponentially in their number.
Definition 15 (Uninformed link).
Given a layer assignment $S_{1},\ldots,S_{d^{\prime}}$, a link $\ell-1\to\ell$ (for $\ell\in\{2,\ldots,k\}$) is informed if position $\ell-1$ is assigned to a strictly earlier layer than position $\ell$, and uninformed otherwise (i.e., $v_{\ell-1}$ is not yet determined when $v_{\ell}$ is committed).
Observation 2 (Uninformed-link count).
Any assignment of $k$ positions into $d^{\prime}<k$ layers leaves at least $k-d^{\prime}$ uninformed links.
Definition 16 (Conditionally $\gamma$-spread distribution).
A distribution $\mathcal{D}$ over $k$-position constraint chains is conditionally $\gamma$-spread if for every $\ell\in\{2,\ldots,k\}$, every $a,b\in[m]$, and every fixing of all constraint functions except the row $P_{\ell}(a,\cdot)$: $\Pr_{\mathcal{D}}[\,b\in P_{\ell}(a,\cdot)\mid\text{all other constraint data}\,]\leq\gamma$.
I.e., membership of any single element in a row has probability $\leq\gamma$, conditioned on the rest of the chain.
Definition 17 (Random constraint distribution).
In the random $(k,m,s)$-distribution, every row ($P_{1}$ and each $P_{\ell}(a,\cdot)$) is an independent uniformly random $s$-subset of $[m]$. This distribution is conditionally $(s/m)$-spread.
Over a distribution on inputs, a strategy may resolve some constraint chains but not others; we measure the probability of resolution.
Theorem 1 (Exponential depth–width separation).
Let $\mathcal{D}$ be a conditionally $\gamma$-spread distribution over $k$-position constraint chains with $\gamma<1$.
  1. (a)
    A sequential strategy with $k$ layers and width $1$ resolves $\mathsf{Spec}_{\mathsf{cg}}$ with probability $1$.
  2. (b)
    For any $d^{\prime}<k$, any deterministic $d^{\prime}$-layer strategy with width $w$ resolves $\mathsf{Spec}_{\mathsf{cg}}$ with probability at most $w\cdot\gamma^{\,k-d^{\prime}}$.
  3. (c)
    Consequently, achieving constant resolution probability with $d^{\prime}<k$ layers requires width $w\geq(1/\gamma)^{k-d^{\prime}}$.
For the random $(k,m,s)$-distribution, $\gamma=s/m$, giving width $w\geq(m/s)^{k-d^{\prime}}$.
Part (a) is Observation 1. Part (c) follows immediately from part (b), which we prove now:
Proof of Theorem 1(b).
Fix a deterministic $d^{\prime}$-layer strategy with width $w$ and layer assignment $S_{1},\ldots,S_{d^{\prime}}$. By Observation 2, at least $t\geq k-d^{\prime}$ links $\ell_{1},\ldots,\ell_{t}$ are uninformed.
Fix a single candidate tuple $(v_{1},\ldots,v_{k})$ and process the uninformed links in order. At link $\ell_{j}$, condition on all constraint data except the row $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$ and on the outcomes at prior links. Under this conditioning, both $v_{\ell_{j}-1}^{*}$ (from an earlier layer or the conditioning) and $v_{\ell_{j}}$ (from the same or a later layer) are fixed, while the row $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$ remains random by the $\gamma$-spread condition. Hence $\Pr[(v_{\ell_{j}-1}^{*},v_{\ell_{j}})\in P_{\ell_{j}}\mid\text{conditioning}]\leq\gamma$. Multiplying across all $t\geq k-d^{\prime}$ links gives success probability at most $\gamma^{k-d^{\prime}}$ per candidate; a union bound over $w$ candidates gives the result. ∎
Corollary 1 (Hard instances for any fixed strategy).
For every $k,m,s$ with $m\geq 2s$, every $d^{\prime}<k$, and every (possibly randomized) $d^{\prime}$-layer strategy with width $w$, there exists a constraint chain on which the strategy resolves $\mathsf{Spec}_{\mathsf{cg}}$ with probability at most $w\cdot(s/m)^{k-d^{\prime}}$.
Theorem 1(b) bounds the expected resolution probability of any deterministic strategy over the random $(k,m,s)$-distribution by $w\cdot(s/m)^{k-d^{\prime}}$, so for each deterministic strategy some chain achieves this bound. Yao’s minimax principle [YAO77] gives the same guarantee for randomized strategies.
Orthogonality.
Determination depth and computational depth are independent axes: the constrained generation task has determination depth $k$ and computational depth $O(1)$ per step, while any computationally hard function has determination depth 0. Within a single problem, however, the two axes interact: Section 4.1 shows that enriching the basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below (Theorem 3).
Remark 1 (Approximate determination and the exchange rate).
When a strategy tolerates constraint violations, the expected violation count is at least $(k-d^{\prime})(1-\gamma)$ by Theorem 1; for the random $(k,m,s)$-distribution this is tight. The parameter $\gamma$ thus controls the exchange rate between distributional knowledge and sequential depth. For a distribution that is not $\gamma$-spread—one with exploitable correlations—a strategy can achieve per-link success probability $\gamma^{\prime}>\gamma$. At $\gamma^{\prime}=1$ (perfect prediction), no layers are needed; at $\gamma^{\prime}=\gamma$ (no exploitable structure), the full $k$ layers are required.
Remark 2 (Generality beyond autoregressive generation).
The strategy model applies to any procedure operating in sequential rounds of parallel refinement, including diffusion models. A diffusion model with $T$ denoising steps is a $T$-round strategy; if $T<k$, the uninformed-link argument applies and the exponential lower bound holds with $d^{\prime}=T$.
3.3 Distributed Extension and Pointer Chasing
The constrained generation task shares the same constraint-chain structure as $k$-step pointer chasing [NW93, PS84], differing only in sparsity $s$, and plays the same role for determination complexity that pointer chasing plays for communication complexity. The following tradeoff extends the width bound to a distributed setting where communication can substitute for width. Consider a $k$-party communication model in which the constraint chain is distributed: player $\ell$ holds $P_{\ell}$ privately and sends $b_{\ell}$ bits to a central referee, who must output an admissible tuple.
Theorem 2 (Depth–width–communication tradeoff).
For the random $(k,m,s)$-distribution in the $k$-party model, any $d$-round protocol with width $w$ and per-player communication $b_{\ell}$ bits satisfies $\log w+\sum_{\ell\in U}b_{\ell}\geq|U|\cdot\log(m/s)$, where $U$ is the set of uninformed links ($|U|\geq k-d$).
Setting $b_{\ell}=0$ recovers the width bound; setting $w=1$ gives a communication lower bound. Both are tight. Proof in Appendix C.2.
Observation 3 (Pointer chasing as degenerate determination).
At sparsity $s=1$ the specification is functional (determination depth 0); hardness appears only when constraints are distributed ($k$-step pointer chasing, $\Omega(n/k)$-bit tradeoff [NW93]). At $s\geq 2$ the specification is relational (determination depth $k$); the exponential width bound applies even to a centralized machine. The transition from polynomial to exponential occurs exactly at $s=1$ vs. $s\geq 2$—the boundary between functions and relations.
4 Structural Results
The exponential separation shows that determination depth is a meaningful complexity measure with concrete consequences. We now establish structural results that explain why the separation works and what it implies for the general theory. The results in this section concern the interaction between determination depth and computational depth; communication re-enters in the distributed setting (Appendix B). The conservation law (Theorem 3) shows that enriching the commitment basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below—explaining the tightness witnessed by the constrained generation task. The oracle characterization (Proposition 1) shows that under a commutative basis in the online setting, even a strategy with access to an unbounded disclosure oracle cannot reduce the number of layers below $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$. Finally, stable matching (Proposition 2) witnesses the orthogonality of the two axes: computational cost is polynomial regardless of determination depth.
4.1 Conservation of sequential depth
A specification may have internal dependency structure—a chain of commitments where each depends on the previous one’s outcome—even in the offline setting, where forward validity constraints are absent. The following theorem says this structure cannot be eliminated.
Theorem 3 (Conservation law for sequential depth).
Let $\Phi_{0}$ be a constant-depth basis for $\mathsf{Spec}$ (Definition 11). For any basis $\Phi$ and any valid determination over $\Phi$ that resolves $\mathsf{Spec}$ in $d$ layers, let $c_{i}$ be the circuit depth of layer $i$ (where the circuit’s input is the current history, which includes all prior commitment events and their effects on the admissible set). Then $\sum_{i=1}^{d}(1+c_{i})\;\geq\;\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec}).$
A richer basis can reduce the number of layers $d$ by bundling multiple constant-depth commitments into a single commitment, but evaluating that commitment requires circuit depth $c_{i}$ that accounts for the bundled dependencies. The bound says the total sequential depth of commitment evaluation—summing $(1+c_{i})$ over layers—cannot fall below $\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$. The bound is not a counting identity: a richer basis could in principle exploit algebraic cancellations to resolve the specification in fewer total sequential steps than the longest dependency chain, and the theorem’s content is that no such shortcut exists. The constrained generation task (Section 3) witnesses tightness: for every split of the $k$ dependency links, the bound is achieved with equality (Appendix C.2). The proof (Appendix C.1) traces the longest path in the $\Phi_{0}$-dependency DAG: each pair either crosses a layer boundary (contributing to $d$) or forces a circuit path within a layer (contributing to $c_{i}$).
4.2 Oracle characterization
The conservation law bounds the sequential depth of commitment evaluation in both the offline and online settings. What if the strategy itself has unbounded computational power between commitments? In the online setting, this question is non-trivial:
Proposition 1 (Oracle power does not reduce depth).
Under a commutative basis $\Phi$ in the online setting, grant the strategy a free disclosure oracle call at the start of every layer (a function that may inspect the entire history, perform unbounded computation, and return any value). Any determination strategy that resolves $\mathsf{Spec}$ still requires at least $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ layers.
Under a commutative basis, the only constraint on co-applying commitments within a layer is forward validity: their combined effect must preserve nonemptiness of the admissible set at all extensions. Forward validity is a property of the specification, not of the strategy’s knowledge, so no oracle can increase the number of commitments that fit in one layer (Example 1). Commutativity is essential: without it, non-commuting commitments can be sequenced within a single layer, collapsing multiple algebraic layers.
The conservation law and the oracle characterization are complementary: the former bounds computation within layers (commitment evaluation), the latter bounds the number of layers themselves (strategy computation between layers cannot reduce it). In the distributed setting, the oracle characterization serves as a single-agent baseline: Appendix B shows that asynchronous multi-agent systems cannot always achieve this baseline, recovering the Halpern–Moses impossibility of common knowledge as a special case and providing quantitative depth predictions beyond the binary threshold.
4.3 Orthogonality: stable matching
Within a single specification, the conservation law couples determination and computational depth. Across specifications, these are independent: neither subsumes the other. Stable matching witnesses this sharply.
Given preference lists for two disjoint sets of $n$ agents, the specification maps the input to the set of all stable matchings [GS62]. The structural theory of stable matchings is organized around rotations—cyclic reassignments that transform one stable matching into an adjacent one in the lattice of stable matchings [IL86, GI89]. Rotations are partially ordered by precedence ($\rho$ must be applied before $\rho^{\prime}$ can be exposed); the resulting rotation poset encodes the determination structure.
Proposition 2 (Stable matching depth).
For any stable matching instance, determination depth under the rotation basis equals the height of the rotation poset [IL86, GI89]. By the Irving–Leather realization theorem, every finite poset arises as a rotation poset, so every determination depth is achievable.
Finding a stable matching is in P [GS62], yet determination depth can be arbitrarily large: computational cost is polynomial regardless of determination depth. The Gale–Shapley algorithm finds a matching in $O(n^{2})$ steps but is not depth-optimal; a strategy that applies all exposed rotations simultaneously at each layer achieves depth equal to the poset height, recovering the parallel algorithm of Garg [GAR20]. Proof and full development in Appendix D.1.
5 Metacomplexity of Determination Depth
The preceding sections establish determination depth as a complexity measure with concrete separations and structural laws. We now ask: how hard is it to compute the determination depth of a given specification? The answer ranges from NP-hard to PSPACE-complete depending on the setting, with the polynomial hierarchy arising as the exact hierarchy of determination depths.
Proposition 3 (NP-hardness in the offline setting).
Computing determination depth is NP-hard in the offline setting: for decision tree synthesis, determination depth equals minimum tree depth, which is NP-hard to compute [HR76].
Proof in Appendix E. Depth here reflects the tree structure: a depth-$d$ decision tree has $d$ levels of nodes, and each level is one round of variable-test commitments.
In the online setting, an adversarial environment adds universal quantification—the environment fixes some variables, the determiner fixes others—and the metacomplexity captures the full polynomial hierarchy.
Theorem 4 (Determination depth captures the polynomial hierarchy).
Given a Boolean formula $\theta(x_{1},\ldots,x_{n})$ encoded by a polynomial-size circuit, the outcome set is $O=\{0,1\}^{n}$ (all variable assignments) and the initial admissible set is $\mathsf{Spec}(H_{0})=O$. The commitment basis consists of pointwise filters “set $x_{i}=b$”; the environment fixes some variables adversarially, the determiner fixes the rest. After all $n$ variables are fixed, the strategy succeeds iff the resulting assignment satisfies $\theta$.
  1. (i)
    For each fixed $k$, “is determination depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete.
  2. (ii)
    For unbounded $k$ (given as input), the problem is PSPACE-complete.
Proof sketch.
Upper bound. The determiner guesses $k$ rounds to control; the environment controls the remaining nkn-k. Each round, the controlling player fixes one variable. The determiner’s choices are existential quantifiers, the environment’s are adversarial (universal). In the worst case these alternate ($\exists\forall\exists\forall\cdots$), and consecutive same-player rounds collapse, giving at most $2k$ alternations ($\Sigma_{2k}^{P}$); for unbounded $k$, PSPACE.
Hardness. Reduce from $\Sigma_{2k}$-QBF for fixed $k$, or from TQBF for unbounded $k$. Construct a specification with $2k$ variables: at odd rounds the determiner fixes an existential variable, at even rounds the environment fixes a universal variable. The admissible set starts as $O$; each round narrows it by fixing one variable. The determiner has a depth-$k$ strategy iff the corresponding QBF is true. Full proof in Appendix E, where the game is formalized as a QBF instance with variable-fixing commitments. ∎
The polynomial hierarchy is thus precisely the hierarchy of determination depths for circuit-encoded specifications. (Offline, the metacomplexity is NP-hard but does not capture the full hierarchy; see Proposition 3.) Where circuit depth and communication complexity parameterize the PH through computational resources, determination depth parameterizes it through a semantic one—layers of irrevocable choice. The correspondence is not the contribution; it is a calibration. An alternating Turing machine has a fixed quantifier prefix determined by the program; here the determiner optimizes which variables to control and in which order, so the quantifier schedule is itself part of the problem. More fundamentally, the paper’s main results—the exponential separation (Theorem 1) and the conservation law (Theorem 3)—are offline results about depth–width and depth–computation tradeoffs within a single specification, a regime not captured by the quantifier alternation that ATMs model.
6 Conclusion
Complexity theory has measured the cost of computing a uniquely determined answer. This paper introduces determination depth to measure the cost of committing to an answer when many are admissible—a semantic complexity measure orthogonal to computation. The exponential separation (Theorem 1) shows this cost is real, the three-way tradeoff (Theorem 2) shows depth, width, and communication are fungible, the PH characterization (Theorem 4) shows it is well-calibrated against the classical hierarchy. In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge as a special case (Theorem 5), and the conservation law provides quantitative depth–cost predictions beyond the binary threshold. The appendices ground the framework in BSP round complexity, chain-of-thought reasoning, stable matching, extensive-form games, and distributed graph coloring; Appendix F surveys related work in detail. A number of directions remain open:
Distributional vs. worst-case hardness. The exponential separation is distributional: for any single fixed chain, a strategy with full knowledge resolves in one layer. Characterizing which restrictions on the strategy’s knowledge or power yield worst-case hardness is open.
Fault tolerance and randomization. In every deterministic, fault-free multi-party setting we have examined, the specification’s semantic structure fully accounts for the known round complexity via determination depth (Appendices $D.4$). The scope of this correspondence is open: fault-tolerant settings and randomized protocols lie outside the current framework, though the Lovász Local Lemma already shows that randomization can collapse determination depth in some settings (Appendix D.4).
Reversible commitments. The framework assumes commitments are irrevocable. In many settings—autoregressive generation with backtracking, transactional rollback, speculative execution—commitments can be reversed at a cost. Extending the commitment algebra to include inverses (a group rather than a monoid) and characterizing the resulting depth–cost tradeoffs is open.
Provenance of determinations. Classical data provenance explains query answers via a commutative semiring over monotone derivations [GKT07]. Determinations introduce a different algebra: a conditionally commutative monoid of layered commitments. The two structures do not align—semiring provenance applies only after a determination has resolved all ambiguity, and cannot represent how alternative commitments would have led to different explanations. A unifying algebraic theory of determination provenance— tracking how commitments at each layer shape downstream explanations—remains open.
Appendix A Formal Definitions for Determination Depth
This appendix collects the formal definitions deferred from Section 2; no new results are claimed. Fix a commitment basis $\Phi$ and a specification $\mathsf{Spec}$ throughout.
The goal is to characterize the parallelism available in a determination $D(H)$ (the commitment subsequence of a history $H$). In the offline setting, all commitment events are consecutive (no environment events interleave), so the entire determination can be analyzed as a single sequence: adjacent commitments that commute can be parallelized into layers, and the depth of $D(H)$ is the minimum number of such layers. In the online setting, environment events may separate consecutive commitments, and commutation is only meaningful for commitments that share the same history prefix. The layering then applies to each maximal run of consecutive commitments (between successive environment events), and the depth of $D(H)$ is the total number of layers across all such runs. $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ is defined (Definition 6) as the minimum worst-case depth over all strategies that resolve $\mathsf{Spec}$.
A.1 Commuting Layers and Depth
We extend the $\cdot$ notation to multisets: $H\cdot L$ is the history obtained by appending all commitments in $L$ after $H$ (in any order).
Definition 18 (Commuting layer).
A finite multiset $L$ of commitments from a common basis $\Phi$ is a commuting layer for $\mathsf{Spec}$ at $H$ if applying its commitments after $H$ is order-independent: for any two listings $\psi_{1},\ldots,\psi_{t}$ and $\psi^{\prime}_{1},\ldots,\psi^{\prime}_{t}$ of the elements of $L$ (respecting multiplicity),
$\mathsf{Spec}(H\cdot\psi_{1}\cdot\psi_{2}\cdots\psi_{t})\;=\;\mathsf{Spec}(H\cdot\psi^{\prime}_{1}\cdot\psi^{\prime}_{2}\cdots\psi^{\prime}_{t}).$
Pairwise commutation (Definition 2) at $H$ is sufficient; every basis used in this paper satisfies this condition.
Definition 19 (Layering and determination depth).
Let $D(H)=\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ be the determination of a history $H$ over $\Phi$. A run is a maximal subsequence of consecutive commitment events in $H$ (with no intervening environment events). A layering of a run $R=\psi_{1}\cdot\cdots\cdot\psi_{r}$ is a partition into nonempty multisets $L_{1},\ldots,L_{k}$ such that each $L_{i}$ is a commuting layer for $\mathsf{Spec}$ at the history $H_{0}\cdot L_{1}\cdots L_{i-1}$, where $H_{0}$ is the history prefix immediately before the run. The depth of a run is the minimum $k$ over all its layerings. The depth of $D(H)$, written $0pt(D(H))$, is the sum of the depths of its runs. In the offline setting, the entire determination is a single run.
For a single run, we write $L_{1}\mathbin{\triangleright}L_{2}\mathbin{\triangleright}\cdots\mathbin{\triangleright}L_{k}$ to denote the layered determination; the $\mathbin{\triangleright}$ notation makes the layering (and hence depth) visible. Every run admits a depth-minimal layering; in the offline setting (a single run), we call this the layered normal form. Depth is bounded by cost ($0pt(D)\leq\mathsf{cost}(D)$, since layers are nonempty), but can be much smaller when most commitments commute.
A.2 Dependency Chains and Depth Characterization
To lower-bound depth, we exhibit commitments that cannot share a layer.
Definition 20 (Forced dependency).
Let $\varphi$ and $\psi$ be commitment events in a determination $D(H)$. We say $\psi$ locally depends on $\varphi$ (in $H$) if, in every layering of $D(H)$, $\varphi$ appears in a strictly earlier layer than $\psi$. We say $\psi$ universally depends on $\varphi$ if this holds in every resolving history whose determination contains both (used only in the distributed setting, Appendix B).
In practice, forced dependency arises when $\psi$ acts as the identity (or violates validity) until $\varphi$ has been applied. The stable matching application (Section D) exhibits this pattern.
Remark 3 (Online depth and forced dependency).
The offline results (Proposition 4 below) use local forced dependency. The distributed results (Theorem 5) use universal forced dependency. In the online setting, depth may exceed the longest dependency chain: run boundaries force additional layers even between independent commitments. The oracle characterization (Proposition 1) handles the online setting directly.
Definition 21 (Dependency chain).
A sequence $\varphi_{1},\varphi_{2},\ldots,\varphi_{k}$ is a dependency chain of length $k$ if each $\varphi_{i+1}$ locally depends on $\varphi_{i}$.
Proposition 4 (Depth equals longest dependency chain (offline)).
In the offline setting, $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ equals the maximum length of a dependency chain over all resolving determinations. In the online setting, the longest chain is a lower bound on depth; run boundaries may force additional layers.
Proof.
A dependency chain of length $k$ requires $k$ layers (each successive commitment must appear in a strictly later layer), so $\mathsf{Depth}_{\Phi}(\mathsf{Spec})\geq$ the longest chain in both settings. In the offline setting (a single run), any determination induces a dependency DAG on its commitments (draw an edge $\varphi\to\psi$ whenever $\psi$ depends on $\varphi$). A depth-minimal layering corresponds to a minimum-height topological layering of this DAG, whose height equals the length of the longest chain. ∎
Appendix B Distributed Determinations and Common Knowledge
This appendix extends the oracle characterization (Section 4.2) to the distributed setting, connecting determination depth to synchronization and common knowledge.
Distributed setting.
We extend the framework to multiple agents. Fix a finite set of agents $\{1,\ldots,n\}$. Each event in a history $H$ is associated with an agent; for agent $p$, the projection $H|_{p}$ is the subgraph of $H$ induced on $p$’s events—local computation steps, message sends, and message receives at pp—with only the edges of $H$ whose both endpoints are events at $p$. (Causal paths through other agents’ events are not visible in the projection; this is the standard Lamport projection.) The partial order on $H|_{p}$ is the transitive closure of $\to$ restricted to $p$’s events: $e_{1}$ precedes $e_{2}$ in $H|_{p}$ whenever $e_{1}\to^{*}e_{2}$ in $H$ and both are events at $p$. The frontier of $H|_{p}$ is the set of sinks of $H|_{p}$ (events with no successors under $\to$).
Indistinguishability.
In the Halpern–Moses framework [HM90], two global states are indistinguishable to agent $p$ if $p$’s local state is the same in both. In our formalism, the analogue of local state is the projection $H|_{p}$. Two global histories $H,H^{\prime}$ are $p$-indistinguishable, written $H\sim_{p}H^{\prime}$, if $H|_{p}$ and $H^{\prime}|_{p}$ are isomorphic as partial orders of typed events—that is, there exists an order-preserving bijection between the events of $H|_{p}$ and $H^{\prime}|_{p}$ that preserves event types (message contents, commitment values, and local computation steps).
Local knowledge and mutual knowledge.
A global property $S$ (a set of histories) is known by agent $p$ at $H$ if $S$ holds at every history $p$-indistinguishable from $H$: $K_{p}(S,H)\;\Leftrightarrow\;\forall H^{\prime}\!:H\sim_{p}H^{\prime}\Rightarrow H^{\prime}\in S$. $S$ is known to everyone at $H$ if $K_{p}(S,H)$ for every $p$; write $E(S,H)$ for this. Define $k$-th order mutual knowledge inductively: $E^{0}(S,H)$ holds iff $H\in S$; $E^{k+1}(S,H)$ holds iff for every agent $p$ and every $H^{\prime}$ with $H\sim_{p}H^{\prime}$, $E^{k}(S,H^{\prime})$ holds. Common knowledge of $S$ at $H$ means $E^{k}(S,H)$ for all $k\geq 0$.
Synchronization points.
We assume each agent’s projection $H|_{p}$ is a chain (a total order on $p$’s events); this holds whenever each agent executes sequentially, as in standard distributed computing models.
[…]
At a synchronization point, every agent sees the same admissible set (condition (iii)), and each agent knows the synchronization occurred (condition (i): $e^{*}_{j,p}$ is in its projection).
Lemma 1 (Synchronization establishes common knowledge).
Assume the synchronization protocol (conditions (i)–(iii)) is common knowledge among all agents. Then at each synchronization point, the current admissible set is common knowledge.
Proof.
Condition (iii) gives $E(S,H)$; common knowledge of the protocol iterates this to all levels. ∎
Connection to the oracle framework.
The oracle characterization (Section 4.2) assumes a single global oracle that sees the entire history. In the distributed setting, each agent has access only to a local oracle that sees its projection $H|_{p}$. A synchronization round (conditions (i)–(iii)) is precisely the mechanism that elevates local oracles to global ones: after synchronization, every agent’s local state includes the same admissible set (by condition (iii) and Lemma 1), so a local oracle can determine the admissible set without global access. Between synchronization points, local oracles are strictly weaker—each sees only its own projection.
Proposition 5 (Determination depth equals synchronization points).
In the online setting under a commutative basis $\Phi$ with $n\geq 2$ agents, if every layer boundary involves a cross-agent forced dependency (Definition 20), then $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ equals the minimum number of synchronization points needed to resolve $\mathsf{Spec}$. When some layers involve only local dependencies, the minimum number of synchronization points may be smaller than $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$, but $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ synchronization points always suffice.
Proof.
Lower bound. Between two consecutive synchronization points, each agent acts on its projection $H|_{p}$ (its local events, received messages, and environment inputs). Within a single layer, an agent can safely apply its own commitments without synchronization: by definition, the commitments in a layer commute and are jointly valid, so each agent’s share can be applied independently. The constraint arises at layer boundaries. By the cross-agent assumption, each layer boundary has a commitment in layer $i+1$ at some agent $q$ that universally depends (Definition 20) on a commitment in layer $i$ at a different agent $p$. Without a synchronization point, $q$ cannot verify that $p$ has completed its layer-$i$ commitment: $q$’s projection may be consistent with histories in which $p$ has not yet acted. Applying a layer-$(i+1)$ commitment in such a history risks invalidity. A synchronization point establishes common knowledge that layer $i$ is complete, enabling all agents to proceed to layer $i+1$. Hence each inter-synchronization interval accomplishes at most one layer, and at least $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ synchronization points are needed.
Upper bound. A synchronous protocol in which all agents exchange projections at each synchronization point can simulate the oracle: the combined projections reconstruct the global history, and the strategy selects the next layer’s commitments. At the synchronization point, every agent knows the global history (from the exchange), knows that every other agent knows it (the exchange was simultaneous), and so on at all levels—establishing common knowledge of the current state. Hence $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ synchronization points suffice. ∎
The following theorem recovers the Halpern–Moses impossibility of asynchronous common knowledge [HM90] as a consequence of the determination framework, via an independent proof that uses only histories, projections, and universal forced dependency.
Theorem 5 (Asynchronous impossibility via determination).
Let $\mathcal{H}_{\mathsf{async}}$ be an asynchronous history class: for every pair of agents $p,q$ and every history $H\in\mathcal{H}_{\mathsf{async}}$ containing an event $e$ at $p$ that is not in the projection of $q$, there exists $H^{\prime}\in\mathcal{H}_{\mathsf{async}}$ with $H\sim_{q}H^{\prime}$ in which $e$ has not occurred. If $\mathsf{Spec}$ has a cross-agent universal forced dependency—a commitment $\varphi$ at some agent $p$ and a commitment $\psi$ at a distinct agent $q$ such that $\psi$ universally depends on $\varphi$ (Definition 20)—then no determination strategy over a commutative basis can resolve $\mathsf{Spec}$ over $\mathcal{H}_{\mathsf{async}}$.
Proof.
Let $\varphi$ at agent $p$ and $\psi$ at agent $q\neq p$ be a cross-agent universal forced dependency. Consider any history $H$ in which $\varphi$ has been applied at $p$. Since $\varphi$ is an event at $p$ and is not in the projection of $q$, the asynchronous condition guarantees a history $H^{\prime}\in\mathcal{H}_{\mathsf{async}}$ with $H\sim_{q}H^{\prime}$ in which $\varphi$ has not occurred. At $H^{\prime}$, $\psi$ is invalid: by universal forced dependency, $\psi$ acts as the identity or violates validity without the prior effect of $\varphi$. Since $q$ cannot distinguish $H$ from $H^{\prime}$, any deterministic strategy that applies $\psi$ at $H$ also applies it at $H^{\prime}$, producing an invalid determination. Hence no strategy can apply both $\varphi$ and $\psi$ without a synchronization point between them, and $\mathsf{Spec}$ is not resolvable over $\mathcal{H}_{\mathsf{async}}$. ∎
Theorem 5—like Halpern-Moses—is a qualitative threshold: synchronization is needed or not. Determination depth refines this to a cost measure: Proposition 5 says at most $k$ synchronization points are needed for depth $k$ (exactly $k$ when every layer boundary involves a cross-agent dependency), and the conservation law (Theorem 3) ensures the total sequential cost cannot be eliminated by enriching the basis. More broadly, the determination framework applies beyond the distributed setting—to single-agent offline problems, combinatorial structures, and autoregressive generation—making the distributed recovery one instantiation of a domain-independent measure.
Corollary 2 (Common knowledge assumptions collapse depth).
If common knowledge of a property $P$ is assumed as part of the model (i.e., $P$ holds at every history in $\mathcal{H}$ and this is common knowledge among all agents), and establishing $P$ through synchronization would require $j$ layers, then the assumption reduces $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ by $j$.
This unifies the diagnostic observations:
  • BSP assumes common knowledge of membership (the process set is fixed and globally known), saving $1$ layer (Section D.4).
  • In extensive-form games, a dominant-strategy mechanism makes every node subgame-trivial, reducing the strategic depth to 0 (Section D.2).
  • LOCAL assumes common knowledge of unique identifiers, collapsing determination depth to $1$ (Section D.4).
In each case, the standard model silently enriches the commitment basis by assuming common knowledge of a property whose establishment would otherwise cost determination layers.
Appendix C Proofs for Conservation and Tightness
C.1 Conservation of sequential depth (Theorem 3)
Let $\Phi_{0}$ be a constant-depth basis for $\mathsf{Spec}$ (Definition 11), and let $G_{\Phi_{0}}$ be its forced-dependency DAG (Definition 20). Each edge $\varphi\to\psi$ in $G_{\Phi_{0}}$ has a concrete witness exclusion: a triple $(H,H\cdot\varphi,o)$ such that excluding outcome $o$ at $H$ is invalid, but excluding $o$ at $H\cdot\varphi$ is valid—the dependency is witnessed by a specific outcome that cannot be excluded until $\varphi$ has taken effect. Any basis $\Phi$ that resolves $\mathsf{Spec}$ must also exclude each witness outcome $o$; the conservation law measures how $\Phi$ pays for these exclusions.
Observation 4 (Semantic dependency implies circuit dependency).
Let $\varphi_{j}$ and $\varphi_{j+1}$ be commitments from a constant-depth basis $\Phi_{0}$ such that $\varphi_{j+1}$ is forced to depend on $\varphi_{j}$. If a circuit $C$ computes a valid layer of a determination that achieves a witness exclusion of both $\varphi_{j}$ and $\varphi_{j+1}$, then $C$ contains a directed path from the gates computing $\varphi_{j}$’s witness exclusion to the gates computing $\varphi_{j+1}$’s witness exclusion.
Proof.
By forced dependency, there exist instances where $\varphi_{j+1}$’s exclusion is invalid unless $\varphi_{j}$’s exclusion has already occurred. If $C$ contained no path from $\varphi_{j}$’s gates to $\varphi_{j+1}$’s, then $\varphi_{j+1}$’s output would be independent of $\varphi_{j}$’s, and on such an instance the circuit would produce an invalid exclusion—a contradiction. ∎
Theorem 3 (Conservation law, restated). For any basis $\Phi$ and any valid determination over $\Phi$ that resolves $\mathsf{Spec}$ in $d$ layers with per-layer circuit depths $c_{1},\ldots,c_{d}$, $\sum_{i=1}^{d}(1+c_{i})\geq d^{*}$, where $d^{*}=\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$ is the determination depth under a constant-depth basis $\Phi_{0}$.
Proof of Theorem 3.
Let $\varphi_{1}\to\cdots\to\varphi_{p}$ be a longest path in $G_{\Phi_{0}}$, so $p=d^{*}$ (Corollary 4). Each edge $\varphi_{j}\to\varphi_{j+1}$ has a witness exclusion $(H_{j},H_{j}\cdot\varphi_{j},o_{j})$: outcome $o_{j}$ cannot be excluded at $H_{j}$ until $\varphi_{j}$ has taken effect. Fix any determination $D$ under basis $\Phi$ with $d$ layers and per-layer circuit depths $c_{1},\ldots,c_{d}$. Since $D$ resolves $\mathsf{Spec}$, each $o_{j}$ is eventually excluded; let $\lambda(j)$ be the $\Phi$-layer in which $o_{j}$ is first excluded. The richer basis $\Phi$ may exclude multiple witness outcomes within a single layer.
For each consecutive pair $\varphi_{j}\to\varphi_{j+1}$ in the path ($j=1,\ldots,p\!-\!1$), exactly one of the following holds.
Case 1: $\lambda(j)<\lambda(j+1)$. The predecessor’s effect is achieved in a strictly earlier $\Phi$-layer. This pair contributes at least $1$ to $d$.
Case 2: $\lambda(j)=\lambda(j+1)=i$. Both effects are achieved in the same $\Phi$-layer $i$. By Observation 4, the circuit computing $\Phi$-layer $i$’s commitments contains a directed path from the gates for $\varphi_{j}$’s exclusion to those for $\varphi_{j+1}$’s, contributing at least $1$ to $c_{i}$.
Case 3: $\lambda(j)>\lambda(j+1)$. The successor’s effect is achieved in an earlier $\Phi$-layer than the predecessor’s. Then $\varphi_{j+1}$’s exclusion is valid without $\varphi_{j}$’s prior exclusion, contradicting forced dependency. Hence Case 3 does not arise.
Each of the p1p-1 consecutive pairs contributes at least $1$ to either $d$ (Case 1: a layer boundary) or some $c_{i}$ (Case 2: a circuit path within a layer), and each layer visited by the path contributes its $1$ term. Hence $\sum_{i=1}^{d}(1+c_{i})\geq p=d^{*}$. ∎
Constrained generation instantiation.
The conservation law applies directly to the constrained generation task (Section 3.1). The outcome space is a product $O=[m]^{k}$, and the constant-depth basis $\Phi_{0}$ is the coordinate basis: each commitment “fix $v_{\ell}=v$” selects a value for one position, which is computable in $O(1)$ circuit depth. The dependency DAG $G_{\Phi_{0}}$ is the chain $1\to 2\to\cdots\to k$, since the feasible values at position $\ell$ depend on the value chosen at position $\ell-1$ (via the successor set $P_{\ell}(v_{\ell-1},\cdot)$). In Case 2 of the conservation-law proof, this dependency becomes a literal circuit data path: the sub-circuit computing $v_{\ell_{j+1}}$ must have $v_{\ell_{j}}$ on its input path, because the successor sets $P_{\ell_{j+1}}(a,\cdot)$ differ across predecessor values $a$. Case 3 cannot arise for the same reason: $v_{\ell_{j+1}}$’s feasibility depends on $v_{\ell_{j}}$’s value, so the successor’s exclusion cannot be valid before the predecessor’s.
The constrained generation task is offline: the constraint chain is fully given, so there is one complete history with no extensions. This rules out branch-specific conditioning—the strategy cannot make different choices for different futures, because there are no branches. The following remark shows why this matters.
Remark 4 (Branch-specific commitments can bypass conservation).
In the online setting, a basis that allows branch-specific commitments—whose effects depend on which future history materializes—can bypass the conservation bound. In the three-valued consensus of Example 1 (an online specification), $\mathsf{Depth}_{\Phi}(\mathsf{Spec})=2$ (under the atomic basis), but a single non-pointwise commitment can choose differently per branch (keep $b$ at $H_{1}$, $c$ at $H_{2}$, $c$ at $H_{3}$), resolving the specification in one layer with $O(1)$ circuit depth.
The offline product-space setting of the constrained generation task rules out this escape: all commitments are applied after the history is fixed, and the coordinate basis is commutative, so no branch-specific conditioning is possible.
C.2 Tightness of the conservation law and the three-way tradeoff
We first prove the three-way tradeoff, then the conservation tightness.
Theorem 2 (Depth–width–communication tradeoff, restated). For the random $(k,m,s)$-distribution in the $k$-party model, any $d$-round protocol with width $w$ and per-player communication $b_{\ell}$ bits satisfies $\log w+\sum_{\ell\in U}b_{\ell}\geq|U|\cdot\log(m/s)$, where $U$ is the set of uninformed links ($|U|\geq k-d$).
Proof of Theorem 2.
Fix a $d$-round protocol with width $w$ and layer assignment $S_{1},\ldots,S_{d}$. Let $U$ be the set of uninformed links. Index the uninformed links as $\ell_{1},\ldots,\ell_{t}$; Observation 2 tells us $t\geq k-d$ .
Message-decoding lemma. Let $R\subseteq[m]$ be a uniformly random $s$-element subset, let $M=M(R)$ be a $b$-bit message (a deterministic function of $R$), and let $g$ be any decoder that outputs an element $g(M)\in[m]$. Then $\Pr[g(M)\in R]\leq\min(1,\;2^{b}\cdot s/m)$. Proof: Count pairs $(R,\mu)$ with $\mu=M(R)$ and $g(\mu)\in R$. For each of the at most $2^{b}$ message values $\mu$, the decoder outputs a fixed element $g(\mu)$. The number of $s$-subsets containing $g(\mu)$ is $\binom{m-1}{s-1}$. Hence the total number of good pairs is at most $2^{b}\cdot\binom{m-1}{s-1}$. Dividing by the total $\binom{m}{s}$ subsets gives $\Pr[g(M)\in R]\leq 2^{b}\cdot\binom{m-1}{s-1}/\binom{m}{s}=2^{b}\cdot s/m$. Taking the minimum with $1$ gives the claim.
Per-link bound. Process the uninformed links in order $\ell_{1},\ldots,\ell_{t}$. At step $j$, condition on all constraint functions $(P_{1},\ldots,P_{k})$ except the row $P_{\ell_{j}}(v_{\ell_{j}-1},\cdot)$, on the outcomes at links $\ell_{1},\ldots,\ell_{j-1}$, and on all messages from players other than $\ell_{j}$. Under this conditioning, $v_{\ell_{j}-1}$ is fixed (from earlier layers or the same layer), $v_{\ell_{j}}$ is a deterministic function of player $\ell_{j}$’s message, and $P_{\ell_{j}}(v_{\ell_{j}-1},\cdot)$ remains a uniformly random $s$-subset (by Definition 17). Applying the message-decoding lemma with $b=b_{\ell_{j}}$ (where $v_{\ell_{j}}$ is determined by player $\ell_{j}$’s message and the fixed side information under the current conditioning), the conditional success probability at link $\ell_{j}$ is at most $\min(1,\;2^{b_{\ell_{j}}}\cdot s/m)$.
Combining. By the chain rule across all uninformed links, the success probability of a single candidate is at most $\prod_{\ell\in U}\min(1,\;2^{b_{\ell}}\cdot s/m)$. A union bound over $w$ candidates gives total success probability at most $w\cdot\prod_{\ell\in U}\min(1,\;2^{b_{\ell}}\cdot s/m)$. For the strategy to succeed with positive probability, this upper bound must be at least $1$. Taking logarithms:
$\log w+\sum_{\ell\in U}\min(0,\;b_{\ell}-\log(m/s))\;\geq\;0.$
Since $\min(0,x)\leq x$,
$0\;\leq\;\log w+\sum_{\ell\in U}\min(0,\;b_{\ell}-\log(m/s))\;\leq\;\log w+\sum_{\ell\in U}(b_{\ell}-\log(m/s)),$
which rearranges to $\log w+\sum_{\ell\in U}b_{\ell}\geq|U|\cdot\log(m/s)$. ∎
Theorem 6 (Tightness of the conservation law).
For the $k$-position constrained generation task (Definition 13) with $m\geq 2s$, every determination using $d$ layers satisfies $\sum_{i=1}^{d}(1+c_{i})\geq k$, and this bound is achieved with equality for every $d\in\{1,\ldots,k\}$.
Proof.
Lower bound. The dependency DAG for the constrained generation task is the chain $1\to 2\to\cdots\to k$ (each position’s feasibility depends on the previous value via $P_{\ell}(v_{\ell-1},\cdot)$), which has longest path $k$. By Theorem 3, $\sum_{i=1}^{d}(1+c_{i})\geq k$ for any $d$-layer determination.
Upper bound. For any $d\in\{1,\ldots,k\}$, partition the $k$ levels into $d$ contiguous blocks $B_{1},\ldots,B_{d}$ of sizes $\lfloor k/d\rfloor$ or $\lceil k/d\rceil$. In layer $i$, commit to all positions in $B_{i}$ by computing them sequentially: given the value $v_{\ell-1}$ from the previous position (either from an earlier layer or from earlier in the same layer’s computation), look up any $v_{\ell}\in P_{\ell}(v_{\ell-1},\cdot)$. Each lookup has circuit depth $O(1)$ (it is a table scan of $P_{\ell}$), and the $|B_{i}|$ lookups within layer $i$ are chained, giving $c_{i}=|B_{i}|-1$. The total is $\sum_{i=1}^{d}(1+(|B_{i}|-1))=\sum_{i=1}^{d}|B_{i}|=k$. ∎
Appendix D Applications to Classical Problems
This section demonstrates the breadth of the determination framework by applying it to problems across several domains: stable matching, extensive-form games, chain-of-thought reasoning, and distributed round complexity (BSP and LOCAL). In each case, the framework either recovers a known result with a new explanation or reveals structure that existing models do not capture. Determination depth arises for two reasons across these examples: enablement, where each commitment creates the sub-problem for the next layer (constrained generation, extensive-form games); and interference, where commitments that are individually valid conflict when applied in the same layer (stable matching under the rotation basis). Each example connects to known results in its domain: stable matching recovers Garg’s parallel algorithm [GAR20] and witnesses the orthogonality of determination and computational depth (finding a stable matching is in P, yet determination depth can be arbitrarily large); extensive-form games measure the game’s strategic depth—the moves where a player must break a tie among equally-optimal options—connecting to Selten’s trembling-hand refinement [SEL75]; chain-of-thought reasoning decomposes CoT length into determination depth, computational depth, and architectural overhead, explaining why longer chains sometimes degrade [SMA+25, SZW+25]; and BSP/LOCAL expose hidden modeling assumptions (Ameloot et al.’s non-obliviousness [ANd13]) that silently collapse determination layers.
The examples progress from a purely offline setting (stable matching) through online single-agent generation (chain-of-thought) to adversarial online interaction (extensive-form games) and distributed multi-agent computation (BSP and LOCAL), illustrating how determination depth provides exact diagnostics across increasing environmental complexity.
D.1 Stable Matching and Determination Universality
Stable matching is a canonical relational specification: given preference lists for two disjoint sets of $n$ agents (encoded as environment events in an offline history), the specification $\mathsf{Spec}_{\mathsf{SM}}$ maps each history to the set of all stable matchings of the encoded instance [GS62]. The number of stable matchings ranges from $1$ to exponentially many, and the specification is inherently relational whenever more than one exists. The analysis below is in the offline setting: all preference lists are given and the specification has no extensions. We show that determination depth for stable matching is exactly characterized by a classical combinatorial invariant—the height of the rotation poset—and that stable matching is universal for determination depth: every finite depth arises as the rotation-poset height of some stable matching instance.
D.1.1 Rotations as commitments
The structural theory of stable matchings is organized around rotations [IL86, GI89]. A rotation is a cyclic reassignment of partners that transforms one stable matching into an adjacent one in the lattice of stable matchings. Formally, a rotation $\rho=(a_{0},b_{0}),(a_{1},b_{1}),\ldots,(a_{r-1},b_{r-1})$ is a sequence of matched pairs in some stable matching such that $b_{i+1\bmod r}$ is the next partner on $a_{i}$’s preference list (after $b_{i}$) with whom $a_{i}$ appears in some stable matching (this can be computed without enumerating all stable matchings [IL86]); applying $\rho$ reassigns each $a_{i}$ to $b_{i+1\bmod r}$, producing an adjacent stable matching in the lattice.
The set of rotations, partially ordered by precedence ($\rho\prec\rho^{\prime}$ if $\rho$ must be applied before $\rho^{\prime}$ can be exposed), forms the rotation poset $\Pi$. The key structural facts (Irving and Leather [IL86], Gusfield and Irving [GI89]) are:
  1. (i)
    The downsets (closed subsets) of $\Pi$ are in bijection with the stable matchings of the instance.
  2. (ii)
    Every finite poset is realizable as the rotation poset of some stable matching instance.
We take rotations as the commitment basis $\Phi_{\mathsf{rot}}$. The commitment for rotation $\rho$ is ambiguity-sensitive (non-pointwise): whether $\rho$ can be applied depends on the current admissible set, not just on individual outcomes. Writing $S=\mathsf{Spec}(H)$ for the admissible set at history $H$:
[…]
where $\mathsf{ds}(\mu)$ denotes the downset of $\Pi$ corresponding to matching $\mu$ (by fact (i) above, each stable matching corresponds to a unique downset). The commitment is irrevocable and satisfies shrinkage (it filters the admissible set rather than transforming matchings; in the determination framework, “applying a rotation” means retaining only matchings consistent with that rotation having been applied). Any determination that applies rotations in a valid topological order preserves feasibility.
D.1.2 Determination depth equals rotation poset height
Proposition 6 (Determination depth of stable matching).
For any stable matching instance with rotation poset $\Pi$,
$\mathsf{Depth}_{\Phi_{\mathsf{rot}}}(\mathsf{Spec}_{\mathsf{SM}})\;=\;\mathrm{height}(\Pi),$
where $\mathrm{height}(\Pi)$ is the length of the longest chain in $\Pi$.
Proof.
Upper bound. Partition the rotations of $\Pi$ into layers by a longest-path layering: layer $i$ contains all rotations whose longest chain of predecessors has length $i$. After layers $1,\ldots,i\!-\!1$ have been applied, every rotation in layer $i$ is exposed: by the downset–matching bijection (i), applying a rotation $\rho^{\prime}$ retains only matchings whose downsets contain $\rho^{\prime}$, so after all predecessors of a layer-$i$ rotation have been applied, every surviving matching’s downset contains them. Within layer $i$, the rotations are pairwise incomparable in $\Pi$; since all are simultaneously exposed, applying any subset does not affect the exposure status of the others. (Exposure of $\rho^{\prime}$ requires all predecessors of $\rho^{\prime}$ to be in $\mathsf{ds}(\mu)$ for every surviving $\mu\in S$. Applying an incomparable $\rho$ only shrinks $S$, which can only make this condition easier to satisfy, not harder.) Hence the commitments within each layer commute, and the number of layers equals $\mathrm{height}(\Pi)$.
Lower bound. Let $\rho_{1}\prec\rho_{2}\prec\cdots\prec\rho_{h}$ be a longest chain in $\Pi$. Each $\varphi_{\rho_{i+1}}$ depends on $\varphi_{\rho_{i}}$ (Definition 20): before $\rho_{i}$ is applied, the admissible set contains matchings whose downsets do not include $\rho_{i}$, so $\rho_{i+1}$ is not exposed and $\varphi_{\rho_{i+1}}$ acts as the identity. This gives a dependency chain of length $h$, so $\mathsf{Depth}_{\Phi_{\mathsf{rot}}}(\mathsf{Spec}_{\mathsf{SM}})\geq h$ by Proposition 4. ∎
Remark 5 (Universality).
By the Irving–Leather realization theorem [GI89], every finite poset arises as a rotation poset. Since determination depth equals poset height (Proposition 6), stable matching realizes every possible determination depth: for every $k$, there exists an instance with determination depth exactly $k$. In particular, finding a stable matching is in P, yet the determination depth can be arbitrarily large—computational cost is polynomial regardless of determination depth.
Depth-optimality of strategies.
The Gale–Shapley algorithm [GS62] finds a stable matching in $O(n^{2})$ sequential steps, but it is not depth-optimal: it uses $O(n^{2})$ sequential steps even when the rotation poset has small height. A depth-optimal strategy—applying all exposed rotations simultaneously at each layer—achieves depth equal to the poset height, potentially much less than $O(n^{2})$. This is precisely the parallel algorithm derived independently by Garg’s lattice-linear predicate framework [GAR20], which the determination framework recovers as a consequence of the depth characterization.
Whether the conservation law yields new lower bounds on parallel stable matching algorithms is an open question.
D.2 Strategic Depth in Extensive-Form Games
How much arbitrary choice does a game force upon a player? Game-tree depth overcounts: at forced moves, where only one option is optimal, the player faces a purely computational burden, not a relational one. We use determination depth to measure a game’s strategic depth: the number of moves where the player must break a tie among multiple equally-optimal options—an irreducible relational cost that no amount of computation can eliminate.
Specification.
Consider a two-player extensive-form game with perfect information. Player 1 (the determiner) and Player 2 (the environment) alternate moves in a game tree $T$; this is an online specification, since player 2’s moves are environment events that extend the history between player 1’s commitments. The outcome set $O$ is the set of all root-to-leaf paths (complete plays), and the initial admissible set is $\mathsf{Spec}(H_{0})=O$. We restrict the admissible set to plays consistent with subgame-perfect equilibrium (SPE) [SEL65]: at each node, the acting player’s move must be optimal given optimal play in all subsequent subgames. Formally, the admissible set at history $H$ is the set of all complete plays extending $H$ that arise under some subgame-perfect equilibrium of the full game. The specification is relational whenever multiple SPE plays pass through the current history.
Commitment basis.
At each player-1 node $v$, the commitment “choose child cc” excludes all plays not passing through $c$: a pointwise filter. Player-2 moves are environment events—outside the determiner’s control, regardless of player-2’s equilibrium rationality. Player-1 moves at subgame-trivial nodes (where the SPE prescribes a unique move) are also effectively environment events: the move is uniquely determined, so no choice is involved. Only moves at subgame-non-trivial nodes—where multiple children lead to plays with the same SPE value for player 1—are genuine commitments requiring an arbitrary choice.
Definition 22 (Subgame-non-trivial node).
A player-1 node $v$ is subgame-non-trivial if at least two children of $v$ are each consistent with some (possibly different) subgame-perfect equilibrium of the full game. It is subgame-trivial if exactly one child is SPE-consistent.
Proposition 7 (Strategic depth of extensive-form games).
For a two-player extensive-form game with perfect information, the determination depth from player 1’s perspective (under the SPE-restricted specification) equals the maximum number of subgame-non-trivial player-1 nodes on any root-to-leaf path.
Proof.
Upper bound. A strategy that auto-plays the unique SPE move at subgame-trivial nodes and commits at subgame-non-trivial nodes uses one layer per non-trivial node along any realized play. Forced moves (at trivial nodes) and player-2 responses occur between commitment layers as environment events, contributing no determination depth.
Lower bound. Consider two consecutive subgame-non-trivial player-1 nodes $u$ and $v$ on a root-to-leaf path, with $v$ deeper than $u$. Node $v$ has SPE-consistent children $c_{1},c_{2}$, so there exist SPEs $\sigma_{1},\sigma_{2}$ of the full game inducing plays through $c_{1}$ and $c_{2}$ respectively. Between $u$ and $v$, player-2 moves and trivial player-1 moves occur as environment events. The SPEs $\sigma_{1}$ and $\sigma_{2}$ may prescribe different player-2 moves at nodes between $u$ and $v$, so the admissible set at vv—which plays remain SPE-consistent—depends on which environment events materialize after the commitment at $u$. Under one sequence of player-2 moves, only plays through $c_{1}$ may remain admissible at $v$; under another, only plays through $c_{2}$. Since these environment events have not occurred when the commitment at $u$ is made, no commitment at $u$ can validly determine the choice at $v$: any fixed choice risks selecting a child whose plays are inadmissible under some SPE-consistent continuation. The choice at $v$ is therefore a forced dependency on the history up to $v$. Chaining these forced dependencies across all non-trivial nodes on the path gives a dependency chain of length equal to their count (Proposition 4). ∎
The result decomposes game-tree depth into two components: strategic depth (the non-trivial nodes, where the player must break a tie among equally-optimal moves) and forced depth (the trivial nodes, where the optimal move is uniquely determined and no choice is involved). This decomposition is distinct from classical game-tree depth and alternating Turing machine quantifier depth, both of which count all player-1 moves—including forced ones that involve no relational choice. The distinction is analogous to the BSP diagnostic (Section D.4): just as BSP overcharges by counting communication rounds that resolve no ambiguity, game-tree depth overcharges by counting moves that present no relational choice. The analysis is restricted to finite perfect-information games (Zermelo’s setting); extending to imperfect information requires handling information sets and is left open.
Remark 6 (Error amplification under bounded rationality).
Strategic depth has consequences beyond the relational burden itself. If a bounded player can compute SPE values but trembles when breaking ties—erring at each non-trivial node with independent probability pp—the probability of perfect play is $(1-p)^{d}$ ($d$ = strategic depth, not game-tree depth $k$). Forced moves present no relational risk; computational failures at forced moves are an orthogonal axis of bounded rationality. Strategic depth is per-player and need not be symmetric: one player can often influence the other’s strategic depth by choosing which subtree to enter. In chess, this is “playing for complications”: steering toward positions where the opponent faces more tie-breaking choices. Balancing payoff maximization against error amplification is an open question that connects to Selten’s trembling-hand refinement [SEL75], with influence over determination depth as a new degree of strategic freedom.
Remark 7 (Mechanism design as basis enrichment).
Beyond extensive-form games, the determination framework applies to mechanism design. A mechanism (e.g., an auction format) can be viewed as enriching the commitment basis: a direct-revelation mechanism allows a player to submit a full strategy as a single commitment, collapsing multiple sequential choices into one layer. A dominant-strategy mechanism (e.g., a second-price auction [VIC61]) makes every node subgame-trivial—each player’s optimal action is independent of others’—collapsing strategic depth to 0. An indirect mechanism (e.g., an ascending auction [MIL00]) trades strategic depth for communication: more rounds of bidding, but less information per round. This echoes the depth–width–communication tradeoff (Theorem 2): formalizing the precise relationship—whether indirect mechanisms obey an analogous conservation law trading determination depth against per-round communication and outcome-set width—is an open problem.
D.3 Decomposing Chain-of-Thought Length
An autoregressive transformer generates tokens sequentially, each an irrevocable commitment. Chain-of-thought (CoT) prompting increases the sequential layers available. Recent work shows CoT provides exponential advantages [FZG+23, LLZ+24, MAA+25], yet longer chains sometimes degrade [SMA+25, SZW+25].
We use the determination framework to decompose CoT length into formally independent components. The decomposition is not a theorem about any specific transformer architecture; it is a semantic lower bound that applies to any autoregressive procedure in which each token is an irrevocable commitment and the task is relational (multiple valid continuations exist). The constrained generation task (Section 3) serves as the witness because it isolates determination depth from computational depth: each commitment is a constant-time table lookup, so the sequential cost is entirely semantic. The total CoT length $T$ is lower-bounded by three formally independent quantities: (i) determination depth dd—non-commuting commitment layers, witnessed by the constrained generation task (Section 3); (ii) computational depth cc—inherent per-step sequentiality from bounded transformer depth (TC0-like [FZG+23]), witnessed by graph connectivity [MAA+25]; and (iii) architectural overhead—tokens spent on information management (context summarization, backtracking, re-derivation), reducible by changing the architecture without changing the task.
Proposition 8 (CoT lower bound from determination depth).
CoT length $T\geq\max(d,c)$: the bound $T\geq d$ follows from Theorem 1 and $T\geq c$ from [FZG+23, MAA+25]. The two bounds are independent: there exist tasks with $d=k$, $c=O(1)$ (Theorem 1), and tasks with $d=0$, $c=\omega(\!\log n)$ [MAA+25]. In both cases, fewer than $\max(d,c)$ layers requires exponential parallel width.
The decomposition explains the apparent conflict. Tasks where CoT provides exponential advantage are determination-bound or computation-bound: their sequential cost is intrinsic, and more layers directly reduce it. Tasks where longer chains degrade [SMA+25, SZW+25] are architecture-bound: $d$ and $c$ are small, so additional tokens are spent on overhead (context management, backtracking, re-derivation) that introduces errors without reducing the bottleneck. For determination-bound tasks, the conditional-spread parameter $\gamma$ governs how much distributional knowledge can substitute for chain length: a model facing a $\gamma$-spread distribution cannot benefit from additional training, while a model facing a more structured distribution can trade prediction quality for fewer layers (Remark 1).
D.4 Diagnosing Distributed Round Complexity
BSP [VAL90] organizes parallel computation into synchronous rounds separated by barriers. We show that BSP round complexity and determination depth disagree in both directions, and that the disagreement is precisely explained by the gap between BSP’s implicit commitment basis and the atomic basis that serves as our intrinsic reference (Section 2.1).
The atomic basis as intrinsic reference.
Under the atomic basis in the online setting, determination depth measures the irreducible commitment structure of a specification—the minimum number of non-commuting layers forced by forward validity constraints, with no assumptions about process identity, membership, or communication primitives. This is the “bare” complexity of the specification. BSP departs from this reference in two ways: it charges for communication (adding rounds where no commitment occurs) and it implicitly enriches the basis (collapsing layers that the atomic basis exposes).
Overcharging: rounds without commitments.
BSP charges one round for any operation requiring communication, whether or not the communication resolves semantic ambiguity. A single-valued specification has determination depth 0 under any basis—the output is uniquely determined, so no commitment is needed. Relational join is the canonical example: parallel implementations that require no synchronization barriers have been known since the early 1990s [WA91], yet BSP assigns one round to the data shuffle because BSP rounds account for communication uniformly. The unnecessary barrier has practical consequences (e.g., straggler sensitivity and inability to pipeline), which the determination framework diagnoses as overcharging for a depth-0 task.
Undercharging: basis enrichment collapses depth.
BSP implicitly enriches the atomic basis by assuming fixed, static membership: the process set is finite, globally known, and does not change during execution. In a real system, establishing membership requires a prior commitment—a discovery protocol, a configuration step, or a leader’s decision about who participates—that BSP treats as given. This assumption hides exactly one unit of determination depth. Under the atomic basis, resolving membership costs at least one layer: a membership commitment (which processes participate?) must precede a value commitment (what does each process output?), and the two do not commute, giving depth $2$. BSP absorbs the membership commitment into its model assumptions, leaving only the value commitment visible—one BSP round for a problem with intrinsic depth $2$. By the conservation law (Theorem 3), the total sequential depth is unchanged; BSP hides the membership layer rather than eliminating it (Corollary 2 in Section B, with $j=1$).
The consequences are sharp. With static membership, universally quantified conditions (e.g., verifying that all processes have completed a round) reduce to finite conjunctions over a known process set [IMM86], and all remaining non-monotonic reasoning can proceed without distributed coordination by the CALM theorem [ANd13]. With dynamic membership, Ameloot’s preconditions fail: universal quantification over participants requires a prior determination of who participates, and each change in membership costs an additional layer.
LOCAL model.
The same diagnostic applies to the LOCAL model of distributed graph computation.
Consider $(\Delta+1)$-coloring of a graph $G=(V,E)$ on $n$ nodes. The specification is relational: many valid colorings exist. The synchronous LOCAL model [LIN92] provides synchronized rounds, simultaneous neighbor-state revelation, and unique node identifiers. Under this basis, $(\Delta+1)$-coloring has determination depth $1$: deterministic symmetry-breaking algorithms (e.g., Cole–Vishkin [CV86]) compute a proper coloring as a function of the ID-labeled neighborhood in $\Theta(\log^{*}n)$ rounds. The round complexity is entirely computational—it measures the cost of symmetry breaking, not of semantic commitment. Common knowledge of unique identifiers collapses the determination structure entirely (Corollary 2).
IDs alone account for the collapse—synchronization and within-layer communication are not needed. Each node waits (asynchronously) for its lower-ID neighbors to commit, then takes the smallest available color. The ID ordering prevents circular dependencies, so no barrier or broadcast is needed. In the language of Ameloot et al. [ANd13], both static membership (Section D.4) and unique identifiers are instances of non-obliviousness: shared knowledge that eliminates the need for distributed coordination. Establishing common knowledge of identifiers lets all remaining non-monotonic reasoning proceed locally, but the conservation law (Theorem 3) ensures that the sequential cost persists as local computation within layers. For coloring, IDs happen to collapse both costs (greedy coloring along the ID-induced orientation is a single-layer local computation), but this is a special property of the coloring specification, not a general consequence of having IDs.
Remark 8 (Randomized strategies).
The analysis above applies to deterministic strategies. With randomization, the Lovász Local Lemma yields $O(1)$-round distributed algorithms for $(\Delta\!+\!1)$-coloring with high probability [MT10], collapsing determination depth to $O(1)$ even without IDs. Randomization thus provides a mechanism for collapsing determination depth by breaking symmetry probabilistically. Characterizing which specifications admit randomized depth collapse is an open question.
Appendix E Metacomplexity Proofs
How hard is it to compute the determination depth of a given specification? This section shows that the answer ranges from NP-hard to PSPACE-complete. In the offline setting, computing determination depth under a pointwise basis is already NP-hard (via decision tree synthesis). In the online setting, the alternation between the determiner’s commitments and the environment’s history extensions produces quantifier alternation: “is depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete for each fixed $k$ and PSPACE-complete for unbounded $k$, so the polynomial hierarchy is precisely the hierarchy of determination depths.
Proposition 9 (Metacomplexity of determination depth).
Computing determination depth is NP-hard in the offline setting: for specifications arising from decision tree synthesis (“given a truth table, output an optimal decision tree”), determination depth equals the minimum decision tree depth, which is NP-hard to compute [HR76].
Proof.
Decision tree synthesis. The environment events encode a truth table $f:\{0,1\}^{n}\to\{0,1\}$ (an offline setting with no further extensions), and the outcome set $O$ is the family of all decision trees on $n$ variables. The specification maps each complete history—encoding a truth table—to the set of decision trees that compute it: $\mathsf{Spec}(H)=\{\,T:T\text{ computes }f_{H}\,\}$, where $f_{H}$ is the truth table encoded in $H$. The commitment basis consists of pointwise filters “test variable $x_{i}$ at the current node”: each such commitment restricts the admissible set to decision trees that test $x_{i}$ at that node. Because each node of a decision tree tests exactly one variable, two test commitments for different variables at the same node have no common tree in their intersection—the admissible set becomes empty—so any valid determination must select exactly one variable per node. A determination of depth $d$ corresponds to a decision tree of depth $d$: each layer selects a variable to test at each node of that layer, and the two test outcomes branch into sub-problems resolved by subsequent layers. Any depth-$d$ determination yields a depth-$d$ tree, and vice versa, so determination depth equals minimum decision tree depth, which is NP-hard to compute [HR76]. Depth here reflects the tree structure: a depth-$d$ decision tree has $d$ levels of nodes, and each level is one round of variable-test commitments. ∎
The NP-hardness result above is an offline result: the specification is fully given and there is no adversarial environment. In the online setting, an adversarial environment adds universal quantification—the environment fixes some variables, the determiner fixes others—and the metacomplexity rises through the full polynomial hierarchy.
To state the complexity result precisely, consider the following class of online specifications. Given a Boolean formula $\theta(x_{1},\ldots,x_{n})$ (encoded by a polynomial-size circuit), define a specification with outcome set $O=\{0,1\}^{n}$ (all variable assignments) and initial admissible set $\mathsf{Spec}(H_{0})=O$ (every assignment is initially admissible). The commitment basis consists of pointwise filters “set variable $x_{i}=b$”: each excludes all assignments with $x_{i}\neq b$, shrinking the admissible set without resolving it completely. The game has $n$ rounds. A level assignment partitions the $n$ rounds between two players: at each determiner round the determiner fixes one variable (an existential choice); at each environment round the environment fixes one variable adversarially (a universal choice). After all $n$ rounds every variable is fixed and the admissible set is a singleton. The determination depth is the number $k$ of determiner rounds; the metacomplexity question is: given $\theta$ and $k$, does there exist a level assignment with $k$ determiner rounds and a strategy for the determiner that guarantees the final assignment satisfies $\theta$, regardless of the environment’s choices? The input size is polynomial in $n$ (the circuit encoding of $\theta$) even though $|O|=2^{n}$.
Theorem 7 (Determination depth captures the polynomial hierarchy).
For the class of specifications above:
  1. (i)
    For each fixed $k$, the problem “is determination depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete.
  2. (ii)
    For unbounded $k$ (given as input), the problem is PSPACE-complete.
This is the standard setup for QBF and PSPACE-complete game evaluation.
Proof.
Upper bound. The determiner guesses which $k$ of the $n$ rounds to control. Each determiner choice is an existential quantifier; each environment choice is a universal quantifier. After all $n$ rounds, every variable is fixed and satisfaction of $\theta$ is checkable in polynomial time by evaluating the circuit. Consecutive rounds controlled by the same player collapse into a single quantifier block, so the quantifier pattern has at most $2k$ alternations, placing the problem in $\Sigma_{2k}^{P}$. When $k$ is part of the input, the number of alternations $2k$ is at most $2n$; since the circuit encoding of $\theta$ has at least $n$ input wires, $2n$ is linear in the input size. A single alternating polynomial-time machine can therefore handle every $k$, placing the problem in APTIME == PSPACE [SIP83].
Hardness. Reduce from $\Sigma_{2k}$-QBF for fixed $k$, or from TQBF for unbounded $k$. Given a quantified Boolean formula
$\exists y_{1}\forall x_{1}\cdots\exists y_{k}\forall x_{k}.\;\theta(x,y)$, construct a specification with $2k$ variables and outcome set $O=\{0,1\}^{2k}$. Odd-numbered rounds ($1,3,\ldots,2k\!-\!1$) are controlled by the determiner, fixing the existential variables $y_{1},\ldots,y_{k}$; even-numbered rounds ($2,4,\ldots,2k$) are controlled by the environment, fixing the universal variables $x_{1},\ldots,x_{k}$. The quantifier alternation of the QBF maps directly onto the round structure: each $\exists$ becomes a determiner round, each $\forall$ an environment round. The initial admissible set is $\mathsf{Spec}(H_{0})=O$ (all assignments); each round’s commitment narrows it by fixing one variable. After all $2k$ rounds the assignment is fully determined; the determiner’s strategy succeeds iff the resulting assignment satisfies $\theta$.
Correctness (the QBF is true iff the determiner has a depth-$k$ strategy). If the QBF is true, the existential player has a winning strategy: a choice of each $y_{i}$ (possibly depending on $x_{1},\ldots,x_{i-1}$) such that $\theta$ is satisfied for every universal assignment. The determiner plays this strategy at the odd rounds, using $k$ layers (one per existential variable), so depth $\leq k$. Conversely, any depth-$k$ determiner strategy is an adaptive assignment to the existential variables that satisfies $\theta$ against every universal response—exactly a witness that the QBF is true. ∎
The polynomial hierarchy is thus precisely the hierarchy of metacomplexity for determination depth: deciding whether a specification has depth $\leq k$ is $\Sigma_{2k}^{P}$-complete, and PSPACE-complete for unbounded $k$. In the offline setting (Proposition 9), the metacomplexity is NP-hard but does not capture the full hierarchy.
Appendix F Related Work
This section positions determination depth among existing notions of staged computation and semantic dependency. Closest antecedents treat staging either as an operational resource (rounds, adaptivity, oracle calls) or as a syntactic discipline (stratification, fixpoint nesting). Our contribution is to identify staging as a semantic invariant of resolution—defined prior to any particular model or language.
Parallelism, depth, and inherent sequentiality.
Depth as a complexity measure has a long history in models of parallel computation. Circuit complexity distinguishes size from depth, isolating irreducible sequential structure even when unbounded parallelism is available [PIP80, BAR90]. Circuit depth measures the inherent sequentiality of evaluating a function: data dependencies force some gates to wait for others. Determination depth measures the inherent sequentiality of choosing an outcome from a relation: commitment dependencies force some decisions to wait for others. The two are orthogonal: a function has determination depth 0 regardless of its circuit depth, while a relation can have high determination depth even when every individual commitment is a constant-depth circuit (Section 3). This orthogonality is witnessed in the offline setting by the constrained generation task (Section 3); in the online setting, forward validity constraints add a further source of determination depth that circuit models do not capture at all. PRAM and BSP models similarly separate local computation from global synchronization, charging depth to barriers or rounds [BC82, VAL90]. In the online (distributed) setting, these models can both undercharge (when a round boundary hides an irrevocable commitment that the model treats as primitive) and overcharge (when communication structure is conflated with semantic resolution); Section D.4 develops this diagnostic in detail. More broadly, classical hierarchy theorems establish strict separations based on time, space, or alternation depth [SIP83]; our determination depth hierarchies are orthogonal, arising from semantic dependency rather than resource bounds.
Trace monoids and partial commutation.
The layered normal form (Section A) resembles the Foata normal form of Mazurkiewicz trace theory [MAZ77]: a canonical factorization of a word in a partially commutative monoid into maximal independent steps. The resemblance is structural but the theories diverge in three ways that produce qualitatively different phenomena. First, in classical trace theory the independence relation is fixed: two letters either commute or they do not, regardless of context. In determination theory, commutation is dynamic: whether two commitments commute depends on the current refined specification, and applying one commitment can create or destroy independence among the remaining ones. This dynamic commutation arises for different reasons in the two settings: in the online setting, forward validity constraints (a commitment that is safe alone may become unsafe after another commitment narrows the admissible set at some extension) create and destroy independence; in the offline setting, enablement (committing to one position determines which choices remain feasible at the next) produces the same effect. Second, this state-dependence makes the analogue of the Foata normal form non-unique—different layering choices lead to different refined specifications and potentially different minimum determination depths—and computing the minimum height becomes NP-hard (Proposition 9), in contrast to the polynomial-time computability of static Foata height. Third, the main results of this paper—the exponential depth–width separation (offline), the oracle characterization (online), and the conservation law (both settings)—have no analogues in classical trace theory. They arise from the semantic content of commitments (feasibility, resolution, forward validity) rather than from the algebraic structure of partial commutation alone. In short, trace theory provides the algebraic skeleton; the semantic content of determination fills it with phenomena that the skeleton alone cannot express.
Oracle models and adaptivity.
The use of a disclosure oracle to characterize depth parallels classical distinctions between adaptive and non-adaptive computation. Decision tree complexity, oracle Turing machines, and communication complexity all exhibit hierarchies based on the number of adaptive rounds [KL80, PS84, NIS91]. Our oracle characterization (Proposition 1) is specific to the online setting under a commutative basis: oracle invocations witness irreducible semantic dependency—points at which the strategy must observe the evolving history before committing further—rather than query adaptivity or information flow. Local computation between oracle calls is unrestricted, emphasizing that determination depth cannot be collapsed by parallelism or control flow alone. In the offline setting, a single oracle call reveals the complete input and the oracle count collapses to one; the sequential cost reappears as computational depth within layers (the conservation law, Theorem 3). In this sense, determination depth and adaptivity depth both measure staged dependency, but the dependency has a different character. An adaptive query reveals information about a fixed, predetermined answer; a determination commitment creates the answer by irrevocably excluding alternatives. Adaptivity depth measures how many times a strategy must look; determination depth measures how many times it must choose. The two are orthogonal (Observation 3, Section 3): pointer chasing at sparsity $s=1$ has high adaptivity depth but determination depth zero (the answer is unique); constrained generation in the offline setting has determination depth $k$ but adaptivity depth zero (the entire input is given).
Round hierarchies in communication complexity.
The closest technical antecedent to our exponential separation is the Nisan–Wigderson round hierarchy for pointer chasing [NW93]; our constrained generation task extends the same chain structure from functions to relations (Observation 3, Theorem 2). The round-elimination technique—reducing a $k$-round protocol to a (k1)(k\!-\!1)-round protocol with a communication blowup—is the standard tool for proving round lower bounds in communication complexity [MNS+98, SEN18]. Mao, Yehudayoff, and Zhang [MYZ25] recently obtained tight pointer-chasing bounds via gadgetless lifting, bypassing round elimination entirely. Both techniques—round elimination and gadgetless lifting—operate in the functional regime ($s=1$), where hardness is informational: the answer is unique but distributed across players. Our lower bound targets the relational regime ($s\geq 2$), where the answer is not unique and hardness arises from the cost of committing to one answer among many, not from distributing information about a fixed answer. Our exponential separation is an offline result (the entire constraint chain is given to a centralized machine), while pointer chasing is an online communication problem (the chain is distributed across players who communicate in rounds). The three-way tradeoff (Theorem 2) unifies both: it interpolates between the offline width bound (setting communication to zero) and the online communication bound (setting width to one). Our lower-bound technique (conditional spread and union bounds over uninformed links) exploits the relational structure of the specification rather than information-theoretic arguments about message content. The key qualitative difference is that round-elimination lower bounds are polynomial in the communication parameter, while our depth–width tradeoff is exponential—and the transition occurs precisely at the boundary between functional and relational specifications ($s=1$ vs. $s\geq 2$).
Logic, stratification, and fixpoints in databases.
Layered semantics appear in database theory through stratified negation, well-founded semantics, and alternating fixpoints [GRS91]; in descriptive complexity, fixpoint nesting and alternation depth classify expressive power [IMM99]. These layerings are properties of a program under a chosen semantics: a stratified Datalog program with $k$ strata requires $k$ non-commuting sealing commitments (one per stratum), but applying well-founded or stable semantics to the same program yields a different commitment structure and potentially a different determination depth. Determination depth, by contrast, is a property of the specification (the mapping from histories to admissible outcome sets), independent of any particular logical formalism or semantics used to define it. Two logic programs that define the same specification have the same determination depth, even if they differ in strata count, fixpoint nesting, or use of negation. Conversely, applying different negation semantics to the same program can yield different specifications and hence different determination depths—a distinction that strata counts alone cannot express.
Leaf languages.
The leaf language framework of Bovet, Crescenzi, and Silvestri [BCS92] characterizes complexity classes by the language-theoretic complexity of the string of outcomes at the leaves of a nondeterministic computation tree: NP corresponds to testing membership in $\{0,1\}^{*}1\{0,1\}^{*}$, PSPACE to a regular language recognizable with $O(\log n)$ memory, and so on. Both frameworks use the structure of a tree of possibilities to classify problems, but the objects and questions differ. Leaf languages classify the acceptance condition applied to a fixed nondeterministic computation; determination depth classifies the commitment structure required to resolve a relational specification. In particular, leaf languages operate on a single computation tree whose branching is fixed by the machine, while determination depth measures layered commitment across histories whose extensions are chosen by an adversarial environment—a distinction that is specific to the online setting. The PH characterization (Theorem 4) recovers the same hierarchy that leaf languages capture, but through alternation of commitment and environment moves in the online game rather than through the complexity of a leaf-string acceptance condition. In the offline setting, the adversarial environment disappears and the exponential separation (Theorem 1) provides a different route to determination depth hierarchies, via commitment dependency rather than alternation. More broadly, the leaf language program provided an elegant classification of existing complexity classes but did not yield new separations or lower bounds. Our framework is aimed at a different target: not reclassifying known classes, but identifying an axis of complexity orthogonal to computation (Section 3), producing new tradeoffs (Theorems 12), and diagnosing existing models (Sections $D.4$).
Coordination and monotonicity.
The Coordination Criterion and the CALM theorem relate monotonicity to the absence of coordination requirements in distributed systems [HEL26, ANd13]. These are inherently online results: coordination cost arises from the need to commit before the full history is known. In our framework, monotone (future-monotone) specifications correspond to the depth-zero fragment, in which no irrevocable commitment is needed at any prefix. Deeper specifications require staged commitments in the online setting, even in the presence of powerful coordination primitives; the distributed extension (Section B) makes this precise.
Inference-time compute and chain-of-thought reasoning.
A growing body of work studies the power and limitations of chain-of-thought reasoning in transformers [FZG+23, LLZ+24, MAA+25]. Feng, Zhang, Gu, Ye, He, and Wang [FZG+23] and Li and Liu [LLZ+24] show that CoT enables bounded-depth transformers (TC0) to solve problems requiring greater computational depth. Mirtaheri, Anil, Agarwal, and Neyshabur [MAA+25] prove an exponential separation between sequential and parallel scaling for graph connectivity. These results concern computational depth—the inherent sequentiality of evaluating a function. Our exponential separation (Section 3) establishes an orthogonal source of sequential advantage: determination depth, arising from non-commuting commitments in relational specifications. The separation is an offline result (the full constraint chain is given); the decomposition of Appendix D.3 extends to the online autoregressive setting, unifying computational depth, determination depth, and architectural overhead into a single framework. Complementary empirical work identifies failure modes of extended reasoning [SMA+25, SZW+25] and studies optimal compute allocation [SLX+24]; our decomposition provides a formal account of why these phenomena arise.
Summary.
Across these domains, notions of depth have appeared as proxies for sequentiality, staging, or adaptivity. Our contribution is to identify determination depth and determination cost as semantic complexity measures—arising from forward validity constraints in the online setting and from commitment dependency in the offline setting—that are orthogonal to computational complexity. Determination depth measures the minimum sequential layers of irrevocable commitment; determination cost measures the total number of commitments; the exponential separation (Theorem 1) shows that reducing one without increasing the other is impossible.
Acknowledgments
Thanks to Peter Alvaro, Tyler Hou, Sarah Morin and Christos Papadimitriou for helpful feedback on earlier drafts.
AI Disclosure.
We used Claude (Anthropic), ChatGPT (OpenAI), and Gemini (Google) to assist with exposition, structural organization, proof review, and adversarial feedback throughout the manuscript. The tools materially affected the prose and presentation in all sections, including section ordering, framing of arguments, and selection of open problems. All mathematical content, definitions, theorems, and proofs are the authors’ own. The authors verified the correctness and originality of all content including references.

### 01 Apr 2026 v2

@@
 Classical complexity theory measures the cost of computing a function, but many computational tasks require committing to one valid output among several.
-We introduce determination depth—the minimum number of sequential layers of irrevocable commitments needed to select a single valid output—and show it is an orthogonal axis measuring the cost of committing.
+We introduce determination depth—the minimum number of sequential layers of irrevocable commitments needed to select a single valid output—and show that no amount of computation can eliminate this cost.
 We exhibit relational tasks whose commitments are constant-time table lookups yet require exponential parallel width to compensate for any reduction in depth.
-A conservation law couples determination and computational depth: enriching the commitment basis trades layers for step complexity, but the total sequential cost is bounded below.
+A conservation law shows that enriching commitments merely relabels determination layers as circuit depth, preserving the total sequential cost.
 For circuit-encoded specifications, the resulting depth hierarchy captures the polynomial hierarchy ($\Sigma_{2k}^{P}$-complete for each fixed $k$, PSPACE-complete for unbounded $k$).
-Determination depth and computational depth are orthogonal: neither subsumes the other.
-Stable matching witnesses the independence sharply—finding a stable matching is in P, yet every finite determination depth arises as the rotation-poset height of some instance.
-In the distributed setting, the framework recovers the Halpern–Moses common-knowledge impossibility.
+In the online setting, determination depth is fully irreducible: unlimited computation between commitment layers cannot reduce their number.
 1 Introduction
@@
 But many tasks require resolving a relation: choosing one admissible outcome from a set, as when a text generator selects one coherent continuation among many, or a distributed system must agree on one valid output among several.
-We show that this distinction creates a new, orthogonal axis of complexity.
-Classical hard functions are computationally expensive but have no cost on this new axis.
-However, there exist relational tasks in which every step is a constant-time table lookup, yet any strategy using fewer than $k$ sequential layers requires exponential width (exponentially many parallel candidates), even with unbounded parallelism within each candidate.
+We show that resolving a relation carries an intrinsic sequential cost that no amount of computation can eliminate.
+Classical hard functions are computationally expensive but have no cost on this axis.
+However, there exist relational tasks in which every step is a constant-time table lookup, yet any strategy using fewer than $k$ sequential layers requires exponential width, even with unbounded parallelism within each layer.
 The bottleneck is not computation, nor communication, but irrevocable commitment.
-This cost is invisible to classical complexity measures—circuit depth, communication complexity, adaptive queries—because they track the cost of computing or distributing an answer, not the cost of committing to one.
+This cost is invisible to classical complexity measures—circuit depth, communication complexity, adaptive queries—which track the cost of computing or distributing an answer, not the cost of committing to one.
 We call this choice process a determination: a collection of irrevocable commitments that narrow the admissible set until a single outcome remains.
@@
 (1) Exponential separation (Section 3): the depth–width tradeoff is exponential, generalizing pointer chasing [NW93, MYZ25] from functions to relations.
-(2) Conservation law (Section 4): enriching the commitment basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below—choices can be bundled into richer steps, but the burden does not disappear.
-The bound is tight.
-Under a commutative basis in the online setting, a supporting oracle characterization shows that determination depth equals the minimum number of observations of the growing history, regardless of computational power between observations.
+(2) Computation cannot help (Section 4): the sequential cost of determination is governed by the dependency structure among commitments.
+Enriching the computational power of commitments merely relabels determination layers as circuit depth within layers—the total sequential cost is conserved (conservation law); the bound is tight.
+Enriching the computational power of the strategy does not help either: under a commutative basis in the online setting, an oracle characterization shows that determination depth equals the minimum number of observations of the growing history, regardless of computation between observations.
 (3) Polynomial-hierarchy characterization (Section 5): for circuit-encoded specifications, “is depth $\leq k$?” is $\Sigma_{2k}^{P}$-complete for each fixed $k$ and PSPACE-complete for unbounded $k$, certifying determination depth as a semantic measure well-calibrated against the classical hierarchy.
@@
 BSP round complexity overcharges by counting communication rounds that involve no commitment, and undercharges by assuming fixed membership.
-In extensive-form games, determination depth measures the game’s strategic depth—the number of moves where a player must break a tie among equally-optimal options—which can be strictly less than the game-tree depth; one player can influence the other’s strategic depth, creating a new dimension of strategic interaction (Appendix D.2).
-In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge [HM90], with the conservation law and exponential separation providing quantitative predictions beyond the binary threshold.
+In extensive-form games, determination depth bounds the game’s strategic depth—the number of moves where a player must break a tie among equally-optimal options—which can be strictly less than the game-tree depth; one player can influence the other’s strategic depth, creating a new dimension of strategic interaction (Appendix D.2).
+In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge [HM90], with the conservation law and exponential separation providing quantitative bounds beyond the binary threshold.
 2 Framework
@@
 We write $H\cdot\varphi$ for the history obtained by adding a commitment event $\varphi$ after all maximal events of $H$ (i.e., $e\to\varphi$ for every $e$ with no successor in $H$), giving $H\sqsubseteq H\cdot\varphi$.
-Adding a commitment can only narrow the admissible set (shrinkage): $\mathsf{Spec}(H\cdot\varphi)\subseteq\mathsf{Spec}(H)$.
+Adding a commitment permanently narrows the admissible set (shrinkage): $\mathsf{Spec}(H^{\prime})\subseteq\mathsf{Spec}(H)$ for every $H^{\prime}\sqsupseteq H\cdot\varphi$.
 In the distributed setting, each agent’s commitments succeed only its own local maximal events, not the global set; Appendix B develops this extension fully.
 Each commitment event $\varphi$ in a history carries its context: the prefix of the history up to $\varphi$ determines the admissible set at which $\varphi$ takes effect.
-Along any chain of extensions, environment events may enlarge or shrink the admissible set, but each commitment event can only shrink it (shrinkage condition).
-A relational specification cannot be resolved to a function by environment events alone; commitments are required, and the cost of those commitments is what determination depth measures (Section 2.1).
+Environment events may change the admissible set in either direction, but commitments can only shrink it, permanently (shrinkage condition).
+Not every relational specification requires commitments; quantifying the requirement is the theme of this work.
 Definition 5 (Determination).
-The determination $D(H)$ of a history $H$ over a commitment basis $\Phi$ is the subsequence $\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ of commitment events in $H$, listed in causal order.
+The determination $D(H)$ of a history $H$ over a commitment basis $\Phi$ is the sequence $\varphi_{1}\cdot\cdots\cdot\varphi_{m}$ ($\varphi_{i}\in\Phi$) of all commitment events in $H$, listed in causal order.
 Its cost is $\mathsf{cost}(D(H))\triangleq m$.
@@
 A determination is valid if each $\varphi_{j}$ preserves nonemptiness: $\mathsf{Spec}(H^{\prime})\neq\emptyset$ for every $H^{\prime}\sqsupseteq H_{j}$.
-A specification is resolved if $|\mathsf{Spec}(H)|=1$ for all $H$; a history resolves $\mathsf{Spec}$ if its determination leaves $|\mathsf{Spec}(H)|=1$.
+The goal is to narrow the admissible set to a single outcome ($|\mathsf{Spec}(H)|=1$); the formal notion of a resolving strategy appears in Definition 6.
 In the offline setting, environment events precede all commitment events, so the determination is a single consecutive block.
@@
 A protocol, algorithm, or agent is a realization of a determination strategy; the complexity results in this paper are stated at the strategy level.
+Example 1 (Consensus server).
+Consensus—agreeing on a single value—is a canonical coordination problem in distributed systems.
+A single sequential server that dictates a choice for the system is a simplified implementation, but sufficient to illustrate the three forces.
+The consensus specification maps a history to a set of candidate values.
+Proposals arrive over time (environment events, extending history $H\sqsubseteq H^{\prime}$), each adding a value to the admissible set at the extended history.
+The strategy’s determination consists of two commitments.
+First, $\varphi_{c}$ closes the proposal phase at some history $H_{c}$: after $H_{c}\cdot\varphi_{c}$, the specification ignores subsequent proposals, so no environment extension of $H_{c}\cdot\varphi_{c}$ can enlarge the admissible set.
+Then $\varphi_{w}$ draws a winner, resolving the specification to one value from the closed proposal set.
+The draw depends on the closure—the strategy cannot choose a winner without knowing the candidates—so the two commitments require two sequential layers.
 Commutation and depth.
@@
 Commutation is relative to a specification and may change as the history grows: $\varphi$ and $\psi$ may commute at $H$ but not at an extension $H^{\prime}\sqsupseteq H$, or vice versa.
-This history-dependence does not arise in classical trace theory, where independence is static, and it is what makes non-commuting commitments impose irreducible sequencing.
-(A history records one linearization of the commitments; commutation ensures the choice of linearization is immaterial.) The depth of a determination is the total number of commuting layers across all runs of consecutive commitments, where commitments within each layer commute pairwise with respect to the specification at that point in the history.
-The full algebraic development appears in Section A.
+This history-dependence does not arise in classical trace theory [MAZ77], where independence is static; in our setting, non-commuting commitments impose irreducible sequencing precisely because their commutation status can shift as the history evolves.
+(A history records one linearization of the commitments; commutation ensures the choice of linearization is immaterial.) The depth of a determination is the minimum number of commuting layers needed to organize its commitments: within each layer, all commitments commute pairwise; across layers, they need not.
+In the online setting, environment events may separate consecutive commitments into distinct runs, and the depth is the total number of layers across all runs (formalized in Section A).
 Definition 6 (Determination cost and determination depth).
@@
 Define $\mathsf{Cost}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;\mathsf{cost}(D(H))$ and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})\triangleq\min_{\sigma}\;\max_{H}\;0pt(D(H))$, where $\sigma$ ranges over strategies that resolve $\mathsf{Spec}$ and $H$ ranges over histories arising under $\sigma$; or $\infty$ if no resolving strategy exists.
-In the offline setting (a single history with no environment extensions), the max is trivial and the definition reduces to a min over determinations.
+In the offline setting (a single complete history), a strategy is simply a determination of that history, and the definition reduces to a min over determinations.
 $\mathsf{Cost}_{\Phi}(\mathsf{Spec})$ is the determination cost and $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ the determination depth.
-Both are properties of the specification, not of any particular strategy—the strategy is the witness, just as a circuit witnesses the depth of a function—and they are the semantic analogues of circuit size and circuit depth: cost counts commitments, depth counts sequential layers.
+Both are properties of the specification, not of any particular strategy—the strategy is the witness, just as a circuit witnesses the depth of a function.
+Cost counts commitments; depth counts sequential layers.
 In the online setting, the max over $H$ gives the environment adversarial control bounded by the history class $\mathcal{H}$; this is a worst-case measure, and the results in this paper focus on this setting.
+In the consensus server example (Example 1), the determination cost is $2$ (close proposals, then draw) and the determination depth is also $2$: the draw depends on the closure, so the two commitments cannot share a layer.
 2.1 Commitment bases
@@
 Atomic commitments are the finest-grained irrevocable commitments: each permanently excludes a single outcome at a single history.
-The permanence condition ($o$ never reappears at any extension) is what makes the exclusion irrevocable; it is satisfied by all specifications studied in this paper.
-In online settings, depth under the atomic basis arises from forward validity constraints: two atomic commitments made at history $H$ cannot be applied in the same layer if their joint effect empties the admissible set at an extension of $H$, as in the following example:
+In online settings, depth under the atomic basis arises from forward validity constraints: two atomic commitments made at history $H$ cannot be applied in the same layer if their joint effect empties the admissible set at an extension of $H$, as in the following consensus example.
+In this variant, environment events may also shrink the admissible set—e.g., by retracting proposals:
 Example 2 (Three-valued consensus).
-Let $O=\{a,b,c\}$ with $\mathsf{Spec}(H_{0})=\{a,b,c\}$, and suppose the environment may extend $H_{0}$ to a history $E\sqsupseteq H_{0}$ with $\mathsf{Spec}(E)=\{a,b\}$.
+Consider a consensus instance (cf.
+Example 1) with three candidate values $O=\{a,b,c\}$ and $\mathsf{Spec}(H_{0})=\{a,b,c\}$.
+Suppose the environment may extend $H_{0}$ to a history $E\sqsupseteq H_{0}$ with $\mathsf{Spec}(E)=\{a,b\}$.
 The refined admissible sets $\mathsf{Spec}(H_{0}\cdot\varphi_{\neg a})$ and $\mathsf{Spec}(E\cdot\varphi_{\neg a})$ are both non-empty ($\{b,c\}$ and $\{b\}$ respectively), so $\varphi_{\neg a}$ is valid at $H_{0}$.
@@
 The intrinsic determination depth of a specification $\mathsf{Spec}$ is $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ where $\Phi$ is the atomic basis.
-Intrinsic depth measures the irreducible layering forced by forward validity constraints alone—the “bare” sequential cost before any computational or architectural assumptions are introduced.
-In the offline setting, forward validity constraints vanish and intrinsic depth collapses; the conservation law (Theorem 3) shows that the sequential cost reappears as computation within layers.
+Intrinsic depth measures the irreducible sequential cost of resolving a specification under the finest-grained commitments, before any computational or architectural assumptions are introduced.
 Properties of the atomic basis.
@@
 Definition 9 (Pointwise basis).
-A commitment $\varphi$ is pointwise if it acts as an outcome-by-outcome filter: whether $o\in\mathsf{Spec}(H\cdot\varphi)$ depends only on $o$ and $H$, not on the rest of $\mathsf{Spec}(H)$.
+A commitment $\varphi$ is pointwise if it acts as an independent filter on each outcome: whether $o\in\mathsf{Spec}(H\cdot\varphi)$ depends only on $o$ and $H$, not on which other outcomes belong to $\mathsf{Spec}(H)$.
 A pointwise basis consists entirely of pointwise commitments.
-A commitment whose effect on one outcome depends on the presence of another is non-pointwise.
 Pointwise commitments always commute (each outcome is filtered independently, so application order is irrelevant).
@@
 The atomic basis is both commutative and constant-depth, making it the canonical minimal basis for the online setting where commitments are made before all information has arrived.
-Under a constant-depth basis, every unit of determination depth corresponds to one unit of sequential time; under a richer basis, some of that time is absorbed into the computation needed to evaluate each commitment.
-The pointwise, commutative, and constant-depth properties each appear in multiple results: pointwise bases arise in the exponential separation (Section 3), the PH characterization (Section 5), and the NP-hardness of computing depth; commutative bases in the oracle characterization (Section 4.2) and the distributed extension (Appendix B); and constant-depth bases in the conservation law (Section 4.1).
+The pointwise, commutative, and constant-depth properties each appear in the results that follow.
 3 Exponential Depth–Width Separation
-We exhibit a family of constrained generation tasks—modeling autoregressive text generation—in which every commitment is a constant-time table lookup, yet any strategy using $d^{\prime}<k$ layers requires width at least $(m/s)^{k-d^{\prime}}$.
+This section proves that determination depth is a meaningful barrier: reducing it requires an exponential blowup in parallel resources, even when each individual commitment is trivial to compute.
+The consensus server (Example 1) required two sequential layers; we now exhibit tasks requiring $k$ layers for any $k$.
+We construct a family of constrained generation tasks—modeling autoregressive text generation—in which every commitment is a constant-time table lookup, yet any strategy using $d^{\prime}<k$ layers requires width at least $(m/s)^{k-d^{\prime}}$.
 3.1 The constrained generation task
 Fix positive integers $k$ (number of positions), $m$ (domain size), and $s$ (constraint sparsity), with $1\leq s\leq m$.
-The task has k1k-1 links $1\to 2\to\cdots\to k$, each constraining a position’s value given its predecessor’s.
 Definition 12 ($k$-position constraint chain).
@@
 $\mathsf{Spec}_{\mathsf{cg}}(P_{1},\ldots,P_{k})\;\triangleq\;\bigl\{\,(v_{1},\ldots,v_{k})\in[m]^{k}\;\bigm|\;v_{1}\in P_{1}\text{ and }(v_{\ell-1},v_{\ell})\in P_{\ell}\text{ for }\ell=2,\ldots,k\,\bigr\}.$
-The specification is relational ($s^{k}$ admissible tuples when $s\geq 2$) and models autoregressive generation: each position is a token, $[m]$ is the vocabulary, and $P_{\ell}$ encodes local coherence constraints.
+The specification is relational ($s^{k}$ admissible tuples when $s\geq 2$) and models autoregressive generation: each position is a token, $[m]$ is the vocabulary, and the k1k-1 links $1\to 2\to\cdots\to k$ encode local coherence constraints.
 Commitment basis.
@@
 Observation 1 (Sequential strategy).
-Committing to one position per layer in order $v_{1},v_{2},\ldots,v_{k}$ always produces a valid tuple (each $v_{\ell}\in P_{\ell}(v_{\ell-1},\cdot)$ exists by sparsity), using $k$ layers and width $1$.
+Committing to one position per layer in order $v_{1},v_{2},\ldots,v_{k}$ always produces a valid tuple—each $v_{\ell}\in P_{\ell}(v_{\ell-1},\cdot)$ exists by sparsity—using $k$ layers and width $1$.
 3.2 Strategy Model and Lower Bound
 We formalize what a strategy with fewer than $k$ layers can do.
-Definition 14 ($d^{\prime}$-layer strategy with width $w$).
-A $d^{\prime}$-layer strategy with width $w$ partitions the $k$ positions into $d^{\prime}$ groups $S_{1},\ldots,S_{d^{\prime}}$ and produces a single history carrying $w$ parallel candidate tuples, each contributing $k$ commitments (one per position) distributed across the $d^{\prime}$ layers.
-In layer $r$, for each candidate, the strategy commits to values at all positions in $S_{r}$; these values may depend on the full input and on earlier-layer commitments, but not on same-layer or later-layer commitments.
-The strategy selects one of the $w$ candidates as its output; it resolves the specification if the output lies in $\mathsf{Spec}_{\mathsf{cg}}$.
-The model is strictly more permissive than autoregressive generation (which fixes left-to-right order and width $1$): it allows arbitrary layer assignments, arbitrary within-layer parallelism, and $w$ independent candidates.
-The only constraint is that commitments within a layer cannot depend on same-layer outcomes—the defining property of parallel execution—so any lower bound here applies to every parallel strategy that respects this constraint, including beam search, speculative decoding, and diffusion-style refinement.
-Each group $S_{r}$ is a commuting layer in which the $w$ candidates each commit to $|S_{r}|$ positions in parallel.
-When $d^{\prime}<k$, some adjacent positions share a layer and cannot observe each other’s values; the lower bound counts these blind spots (uninformed links) and shows that $w$ grows exponentially in their number.
 Definition 14 (Uninformed link).
@@
 Any assignment of $k$ positions into $d^{\prime}<k$ layers leaves at least $k-d^{\prime}$ uninformed links.
+Definition 15 ($d^{\prime}$-layer strategy with width $w$).
+A $d^{\prime}$-layer strategy with width $w$ partitions the $k$ positions into $d^{\prime}$ groups $S_{1},\ldots,S_{d^{\prime}}$ and builds $w$ candidate tuples in parallel.
+In layer $r$, the strategy assigns values to the positions in group $S_{r}$ for each candidate, depending only on the full input and earlier-layer assignments.
+The strategy resolves the specification if any candidate lies in $\mathsf{Spec}_{\mathsf{cg}}$; the final selection among valid candidates adds at most one layer, which does not affect the asymptotic lower bound.
+At each uninformed link, the strategy must guess a value without knowing its predecessor; the theorem below shows that each guess fails with probability controlled by a distributional parameter $\gamma$.
 Definition 16 (Conditionally $\gamma$-spread distribution).
 A distribution $\mathcal{D}$ over $k$-position constraint chains is conditionally $\gamma$-spread if for every $\ell\in\{2,\ldots,k\}$, every $a,b\in[m]$, and every fixing of all constraint functions except the row $P_{\ell}(a,\cdot)$: $\Pr_{\mathcal{D}}[\,b\in P_{\ell}(a,\cdot)\mid\text{all other constraint data}\,]\leq\gamma$.
-I.e., membership of any single element in a row has probability $\leq\gamma$, conditioned on the rest of the chain.
+In short, membership of any one element in a row has probability $\leq\gamma$, conditioned on the rest of the chain.
 Definition 17 (Random constraint distribution).
@@
 For the random $(k,m,s)$-distribution, $\gamma=s/m$, giving width $w\geq(m/s)^{k-d^{\prime}}$.
+The model is strictly more permissive than autoregressive generation (which fixes left-to-right order and width $1$): it allows arbitrary layer assignments, arbitrary within-layer parallelism, and $w$ independent candidates.
+The only constraint is that commitments within a layer cannot depend on same-layer outcomes—the defining property of parallel execution—so the lower bound applies to every parallel strategy that respects this constraint, including beam search, speculative decoding, and diffusion-style refinement.
 Part (a) is Observation 1. Part (c) follows immediately from part (b), which we prove now:
@@
 By Observation 2, at least $t\geq k-d^{\prime}$ links $\ell_{1},\ldots,\ell_{t}$ are uninformed.
-Fix a single candidate tuple $(v_{1},\ldots,v_{k})$ and process the uninformed links in order.
-At link $\ell_{j}$, condition on all constraint data except the row $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$ and on the outcomes at prior links.
-Under this conditioning, both $v_{\ell_{j}-1}^{*}$ (from an earlier layer or the conditioning) and $v_{\ell_{j}}$ (from the same or a later layer) are fixed, while the row $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$ remains random by the $\gamma$-spread condition.
-Hence $\Pr[(v_{\ell_{j}-1}^{*},v_{\ell_{j}})\in P_{\ell_{j}}\mid\text{conditioning}]\leq\gamma$.
+Fix a single candidate tuple $(v_{1},\ldots,v_{k})$ and consider an uninformed link $\ell_{j}$.
+The strategy picks $v_{\ell_{j}}$ without seeing $P_{\ell_{j}}(v_{\ell_{j}-1}^{*},\cdot)$, which is still $\gamma$-spread by assumption, so $\Pr[(v_{\ell_{j}-1}^{*},v_{\ell_{j}})\in P_{\ell_{j}}]\leq\gamma$.
 Multiplying across all $t\geq k-d^{\prime}$ links gives success probability at most $\gamma^{k-d^{\prime}}$ per candidate; a union bound over $w$ candidates gives the result.
@@
 For every $k,m,s$ with $m\geq 2s$, every $d^{\prime}<k$, and every (possibly randomized) $d^{\prime}$-layer strategy with width $w$, there exists a constraint chain on which the strategy resolves $\mathsf{Spec}_{\mathsf{cg}}$ with probability at most $w\cdot(s/m)^{k-d^{\prime}}$.
-Theorem 1(b) bounds the expected resolution probability of any deterministic strategy over the random $(k,m,s)$-distribution by $w\cdot(s/m)^{k-d^{\prime}}$, so for each deterministic strategy some chain achieves this bound.
+Theorem 1(b) bounds the expected resolution probability of any deterministic strategy over the random $(k,m,s)$-distribution by $w\cdot(s/m)^{k-d^{\prime}}$, so for each deterministic strategy some chain exhibits this bound.
 Yao’s minimax principle [YAO77] gives the same guarantee for randomized strategies.
-Orthogonality.
-Determination depth and computational depth are independent axes: the constrained generation task has determination depth $k$ and computational depth $O(1)$ per step, while any computationally hard function has determination depth 0.
-Within a single problem, however, the two axes interact: Section 4.1 shows that enriching the basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below (Theorem 3).
 Remark 1 (Approximate determination and the exchange rate).
 When a strategy tolerates constraint violations, the expected violation count is at least $(k-d^{\prime})(1-\gamma)$ by Theorem 1; for the random $(k,m,s)$-distribution this is tight.
-The parameter $\gamma$ thus controls the exchange rate between distributional knowledge and sequential depth.
-For a distribution that is not $\gamma$-spread—one with exploitable correlations—a strategy can achieve per-link success probability $\gamma^{\prime}>\gamma$.
-At $\gamma^{\prime}=1$ (perfect prediction), no layers are needed; at $\gamma^{\prime}=\gamma$ (no exploitable structure), the full $k$ layers are required.
+The spread parameter $\gamma$ controls the exchange rate between distributional structure and sequential depth: a distribution with more exploitable structure has a smaller effective $\gamma$, reducing the layers needed.
+At one extreme ($\gamma\to 0$, perfect prediction), no layers are needed; at the other ($\gamma=s/m$, no exploitable structure), the full $k$ layers are required.
 Remark 2 (Generality beyond autoregressive generation).
@@
 For the random $(k,m,s)$-distribution in the $k$-party model, any $d$-round protocol with width $w$ and per-player communication $b_{\ell}$ bits satisfies $\log w+\sum_{\ell\in U}b_{\ell}\geq|U|\cdot\log(m/s)$, where $U$ is the set of uninformed links ($|U|\geq k-d$).
-Setting $b_{\ell}=0$ recovers the width bound; setting $w=1$ gives a communication lower bound. Both are tight. Proof in Appendix C.2.
+Setting $b_{\ell}=0$ recovers the width bound; setting $w=1$ gives a communication lower bound. Both are tight. Proof in Appendix C.
 Observation 3 (Pointer chasing as degenerate determination).
-At sparsity $s=1$ the specification is functional (determination depth 0); hardness appears only when constraints are distributed ($k$-step pointer chasing, $\Omega(n/k)$-bit tradeoff [NW93]).
+At sparsity $s=1$ the specification is functional (determination depth 0); the classical round complexity of $k$-step pointer chasing [NW93] arises entirely from distributing the chain across agents, not from relational choice.
 At $s\geq 2$ the specification is relational (determination depth $k$); the exponential width bound applies even to a centralized machine.
-The transition from polynomial to exponential occurs exactly at $s=1$ vs.
+The transition occurs exactly at $s=1$ vs.
 $s\geq 2$—the boundary between functions and relations.
@@
 We now establish structural results that explain why the separation works and what it implies for the general theory.
-The results in this section concern the interaction between determination depth and computational depth; communication re-enters in the distributed setting (Appendix B).
-The conservation law (Theorem 3) shows that enriching the commitment basis trades determination layers for computational depth within layers, but the total sequential cost is bounded below—explaining the tightness witnessed by the constrained generation task.
-The oracle characterization (Proposition 1) shows that under a commutative basis in the online setting, even a strategy with access to an unbounded disclosure oracle cannot reduce the number of layers below $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$.
-Finally, stable matching (Proposition 2) witnesses the orthogonality of the two axes: computational cost is polynomial regardless of determination depth.
+The sequential cost of determination is governed by the dependency structure among commitments—a structure that computation cannot alter.
+The conservation law (Theorem 3) shows that enriching the commitment basis merely relabels determination layers as circuit depth within layers—the total sequential cost is conserved.
+The oracle characterization (Proposition 1) shows that under a commutative basis in the online setting, even unbounded computation between layers cannot reduce their number.
+Communication re-enters in the distributed setting (Appendix B).
 4.1 Conservation of sequential depth
 A specification may have internal dependency structure—a chain of commitments where each depends on the previous one’s outcome—even in the offline setting, where forward validity constraints are absent.
-The following theorem says this structure cannot be eliminated.
+A richer basis can reduce the number of determination layers $d$ by bundling multiple constant-depth commitments into a single commitment, but evaluating that commitment requires circuit depth $c_{i}$ that accounts for the bundled dependencies.
+The total sequential depth of commitment evaluation—summing $(1+c_{i})$ over layers—cannot fall below $\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$.
+This is not a counting identity: a richer basis could in principle exploit algebraic cancellations to resolve the specification in fewer total sequential steps than the longest dependency chain, and the following theorem’s content is that no such shortcut exists.
 Theorem 3 (Conservation law for sequential depth).
@@
 Then $\sum_{i=1}^{d}(1+c_{i})\;\geq\;\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec}).$
-A richer basis can reduce the number of layers $d$ by bundling multiple constant-depth commitments into a single commitment, but evaluating that commitment requires circuit depth $c_{i}$ that accounts for the bundled dependencies.
-The bound says the total sequential depth of commitment evaluation—summing $(1+c_{i})$ over layers—cannot fall below $\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$.
-The bound is not a counting identity: a richer basis could in principle exploit algebraic cancellations to resolve the specification in fewer total sequential steps than the longest dependency chain, and the theorem’s content is that no such shortcut exists.
-The constrained generation task (Section 3) witnesses tightness: for every split of the $k$ dependency links, the bound is achieved with equality (Appendix C.2).
-The proof (Appendix C.1) traces the longest path in the $\Phi_{0}$-dependency DAG: each pair either crosses a layer boundary (contributing to $d$) or forces a circuit path within a layer (contributing to $c_{i}$).
+Each edge $\varphi_{j}\to\varphi_{j+1}$ in the forced-dependency DAG of $\Phi_{0}$ has a witness exclusion: a triple $(H_{j},H_{j}\cdot\varphi_{j},o_{j})$ such that excluding $o_{j}$ at $H_{j}$ is invalid, but excluding $o_{j}$ at $H_{j}\cdot\varphi_{j}$ is valid.
+The proof traces the longest path in this DAG and shows that each edge must be “paid for” by either a layer boundary or a circuit path within a layer.
+The key step is:
+Observation 4 (Semantic dependency implies circuit dependency).
+If $\varphi_{j+1}$ is forced to depend on $\varphi_{j}$ (under a constant-depth basis $\Phi_{0}$), and a circuit $C$ computes a layer that achieves the witness exclusions of both, then $C$ contains a directed path from the gates for $\varphi_{j}$’s exclusion to those for $\varphi_{j+1}$’s.
+Proof.
+By forced dependency, $\varphi_{j+1}$’s exclusion is invalid unless $\varphi_{j}$’s has occurred.
+If $C$ had no such path, $\varphi_{j+1}$’s output would be independent of $\varphi_{j}$’s.
+But the witness exclusion guarantees instances where excluding $o_{j+1}$ without first excluding $o_{j}$ is invalid, so $C$ would produce an invalid refinement on such an instance—a contradiction.
+∎
+Proof of Theorem 3.
+Let $d^{*}=\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$ and let $\varphi_{1}\to\cdots\to\varphi_{d^{*}}$ be a longest forced-dependency path in the $\Phi_{0}$-dependency DAG (Proposition 3).
+Each edge $\varphi_{j}\to\varphi_{j+1}$ has a witness exclusion.
+Fix any determination under basis $\Phi$ with $d$ layers and per-layer circuit depths $c_{1},\ldots,c_{d}$.
+Since the determination resolves $\mathsf{Spec}$, each $o_{j}$ is eventually excluded; let $\lambda(j)$ be the $\Phi$-layer in which $o_{j}$ is first excluded from the admissible set at all extensions.
+For each consecutive pair $\varphi_{j}\to\varphi_{j+1}$, exactly one of the following holds.
+Case 1: $\lambda(j)<\lambda(j+1)$—the pair crosses a layer boundary, contributing at least $1$ to $d$.
+Case 2: $\lambda(j)=\lambda(j+1)=i$—both exclusions occur in the same layer; by forced dependency, the circuit computing layer $i$ must contain a directed path from the gates for $o_{j}$’s exclusion to those for $o_{j+1}$’s, contributing at least $1$ to $c_{i}$.
+Case 3: $\lambda(j)>\lambda(j+1)$—the successor is excluded before the predecessor, contradicting forced dependency.
+Each of the $d^{*}-1$ pairs contributes at least $1$ to either $d$ or some $c_{i}$, and each layer visited contributes its $1$ term. Hence $\sum_{i=1}^{d}(1+c_{i})\geq d^{*}$. ∎
+The constrained generation task (Section 3) witnesses tightness: for every split of the $k$ dependency links, the bound is achieved with equality (Appendix C).
 4.2 Oracle characterization
-The conservation law bounds the sequential depth of commitment evaluation in both the offline and online settings.
-What if the strategy itself has unbounded computational power between commitments?
+The conservation law shows that enriching commitments cannot reduce the total sequential cost.
+What if instead the strategy has unbounded computational power between commitments?
 In the online setting, this question is non-trivial:
 Proposition 1 (Oracle power does not reduce depth).
-Under a commutative basis $\Phi$ in the online setting, grant the strategy a free disclosure oracle call at the start of every layer (a function that may inspect the entire history, perform unbounded computation, and return any value).
+Under a commutative basis $\Phi$ in the online setting, grant the strategy a free call to a disclosure oracle at the start of every layer (a function that may inspect the entire history, perform unbounded computation, and return any value).
 Any determination strategy that resolves $\mathsf{Spec}$ still requires at least $\mathsf{Depth}_{\Phi}(\mathsf{Spec})$ layers.
@@
 Commutativity is essential: without it, non-commuting commitments can be sequenced within a single layer, collapsing multiple algebraic layers.
-The conservation law and the oracle characterization are complementary: the former bounds computation within layers (commitment evaluation), the latter bounds the number of layers themselves (strategy computation between layers cannot reduce it).
+4.3 Relationship to computational complexity
+Determination depth measures a cost that classical complexity does not track.
+A function has determination depth 0 regardless of its circuit depth; a relation can require exponential width to compensate for any reduction in depth, even when every individual commitment is a constant-time operation (Theorem 1).
+This width cost is purely relational: at sparsity $s=1$ (a function), it vanishes entirely (Observation 3).
+In the online setting, the oracle characterization shows that determination depth is irreducible: no amount of computation between layers can reduce their number.
+Classical computational complexity plays a narrow role within the framework: each commitment is itself a function, so its evaluation has classical circuit depth.
+The conservation law shows that this is all computation can do—it emulates atomic-basis determination layers one-for-one as circuit layers within a richer commitment, preserving the total sequential cost exactly.
+Determination depth under the atomic basis is thus a universal lower bound on the total sequential cost of resolution, regardless of the basis used.
+Determination depth on the atomic basis therefore provides a complete accounting of the sequential cost of resolving a relation—both commitment dependency and forward validity, in both offline and online settings.
+Computational depth captures only the cost of evaluating individual commitments, a sub-problem that arises when the basis is enriched and only in settings where forward validity is absent.
 In the distributed setting, the oracle characterization serves as a single-agent baseline: Appendix B shows that asynchronous multi-agent systems cannot always achieve this baseline, recovering the Halpern–Moses impossibility of common knowledge as a special case and providing quantitative depth predictions beyond the binary threshold.
-4.3 Orthogonality: stable matching
-Within a single specification, the conservation law couples determination and computational depth.
-Across specifications, these are independent: neither subsumes the other.
-Stable matching witnesses this sharply.
-Given preference lists for two disjoint sets of $n$ agents, the specification maps the input to the set of all stable matchings [GS62].
-The structural theory of stable matchings is organized around rotations—cyclic reassignments that transform one stable matching into an adjacent one in the lattice of stable matchings [IL86, GI89].
-Rotations are partially ordered by precedence ($\rho$ must be applied before $\rho^{\prime}$ can be exposed); the resulting rotation poset encodes the determination structure.
-Proposition 2 (Stable matching depth).
-For any stable matching instance, determination depth under the rotation basis equals the height of the rotation poset [IL86, GI89].
-By the Irving–Leather realization theorem, every finite poset arises as a rotation poset, so every determination depth is achievable.
-Finding a stable matching is in P [GS62], yet determination depth can be arbitrarily large: computational cost is polynomial regardless of determination depth.
-The Gale–Shapley algorithm finds a matching in $O(n^{2})$ steps but is not depth-optimal; a strategy that applies all exposed rotations simultaneously at each layer achieves depth equal to the poset height, recovering the parallel algorithm of Garg [GAR20].
-Proof and full development in Appendix D.1.
 5 Metacomplexity of Determination Depth
@@
 An alternating Turing machine has a fixed quantifier prefix determined by the program; here the determiner optimizes which variables to control and in which order, so the quantifier schedule is itself part of the problem.
-More fundamentally, the paper’s main results—the exponential separation (Theorem 1) and the conservation law (Theorem 3)—are offline results about depth–width and depth–computation tradeoffs within a single specification, a regime not captured by the quantifier alternation that ATMs model.
+The alternation arises from the interaction between commitments and environment extensions, not from the encoding of the specification.
+More fundamentally, the paper’s main contributions—the exponential separation (Theorem 1), the conservation law (Theorem 3), and the oracle characterization (Proposition 1)—establish that determination depth resists computation in both the offline and online settings, a phenomenon not captured by quantifier alternation.
 6 Conclusion
 Complexity theory has measured the cost of computing a uniquely determined answer.
-This paper introduces determination depth to measure the cost of committing to an answer when many are admissible—a semantic complexity measure orthogonal to computation.
-The exponential separation (Theorem 1) shows this cost is real, the three-way tradeoff (Theorem 2) shows depth, width, and communication are fungible, the PH characterization (Theorem 4) shows it is well-calibrated against the classical hierarchy.
-In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge as a special case (Theorem 5), and the conservation law provides quantitative depth–cost predictions beyond the binary threshold.
+This paper introduces determination depth to measure the cost of committing to an answer when many are admissible—a cost that no amount of computation can eliminate.
+The exponential separation (Theorem 1) shows this cost is real, the conservation law (Theorem 3) and oracle characterization (Proposition 1) show that computation cannot reduce it, and the PH characterization (Theorem 4) shows it is well-calibrated against the classical hierarchy.
+In the distributed setting, the framework recovers the Halpern–Moses impossibility of asynchronous common knowledge as a special case (Appendix B), and the depth–width–communication tradeoff (Theorem 2) shows that the depth, width, and communication costs of resolving a specification are fungible.
 The appendices ground the framework in BSP round complexity, chain-of-thought reasoning, stable matching, extensive-form games, and distributed graph coloring; Appendix F surveys related work in detail.
-A number of directions remain open:
+A number of directions are open:
 Distributional vs.
@@
 Determinations introduce a different algebra: a conditionally commutative monoid of layered commitments.
-The two structures do not align—semiring provenance applies only after a determination has resolved all ambiguity, and cannot represent how alternative commitments would have led to different explanations.
+The two structures do not align naturally.
 A unifying algebraic theory of determination provenance— tracking how commitments at each layer shape downstream explanations—remains open.
@@
 A.1 Commuting Layers and Depth
-We extend the $\cdot$ notation to multisets: $H\cdot L$ is the history obtained by appending all commitments in $L$ after $H$ (in any order).
+We extend the $\cdot$ notation to sets: $H\cdot L$ is the history obtained by appending all commitments in $L$ after $H$ (in any order).
 Definition 18 (Commuting layer).
-A finite multiset $L$ of commitments from a common basis $\Phi$ is a commuting layer for $\mathsf{Spec}$ at $H$ if applying its commitments after $H$ is order-independent: for any two listings $\psi_{1},\ldots,\psi_{t}$ and $\psi^{\prime}_{1},\ldots,\psi^{\prime}_{t}$ of the elements of $L$ (respecting multiplicity),
+A finite set $L$ of commitments from a common basis $\Phi$ is a commuting layer for $\mathsf{Spec}$ at $H$ if applying its commitments consecutively after $H$ (with no intervening environment events) is order-independent: for any two listings $\psi_{1},\ldots,\psi_{t}$ and $\psi^{\prime}_{1},\ldots,\psi^{\prime}_{t}$ of the elements of $L$,
 $\mathsf{Spec}(H\cdot\psi_{1}\cdot\psi_{2}\cdots\psi_{t})\;=\;\mathsf{Spec}(H\cdot\psi^{\prime}_{1}\cdot\psi^{\prime}_{2}\cdots\psi^{\prime}_{t}).$
@@
 A run is a maximal subsequence of consecutive commitment events in $H$ (with no intervening environment events).
-A layering of a run $R=\psi_{1}\cdot\cdots\cdot\psi_{r}$ is a partition into nonempty multisets $L_{1},\ldots,L_{k}$ such that each $L_{i}$ is a commuting layer for $\mathsf{Spec}$ at the history $H_{0}\cdot L_{1}\cdots L_{i-1}$, where $H_{0}$ is the history prefix immediately before the run.
+A layering of a run $R=\psi_{1}\cdot\cdots\cdot\psi_{r}$ is a partition into nonempty sets $L_{1},\ldots,L_{k}$ such that each $L_{i}$ is a commuting layer for $\mathsf{Spec}$ at the history $H_{0}\cdot L_{1}\cdots L_{i-1}$, where $H_{0}$ is the history prefix immediately before the run.
 The depth of a run is the minimum $k$ over all its layerings.
@@
 In the offline setting, the entire determination is a single run.
-For a single run, we write $L_{1}\mathbin{\triangleright}L_{2}\mathbin{\triangleright}\cdots\mathbin{\triangleright}L_{k}$ to denote the layered determination; the $\mathbin{\triangleright}$ notation makes the layering (and hence depth) visible.
 Every run admits a depth-minimal layering; in the offline setting (a single run), we call this the layered normal form.
@@
 The distributed results (Theorem 5) use universal forced dependency.
-In the online setting, depth may exceed the longest dependency chain: run boundaries force additional layers even between independent commitments.
-The oracle characterization (Proposition 1) handles the online setting directly.
+In the online setting, the oracle characterization (Proposition 1) handles depth directly, without reducing to dependency chains.
 Definition 21 (Dependency chain).
@@
 Fix a finite set of agents $\{1,\ldots,n\}$.
-Each event in a history $H$ is associated with an agent; for agent $p$, the projection $H|_{p}$ is the subgraph of $H$ induced on $p$’s events—local computation steps, message sends, and message receives at pp—with only the edges of $H$ whose both endpoints are events at $p$.
-(Causal paths through other agents’ events are not visible in the projection; this is the standard Lamport projection.) The partial order on $H|_{p}$ is the transitive closure of $\to$ restricted to $p$’s events: $e_{1}$ precedes $e_{2}$ in $H|_{p}$ whenever $e_{1}\to^{*}e_{2}$ in $H$ and both are events at $p$.
-The frontier of $H|_{p}$ is the set of sinks of $H|_{p}$ (events with no successors under $\to$).
+Each event in a history $H$ is associated with an agent; for agent $p$, the projection $H|_{p}$ is the restriction of $H$ to $p$’s events—local computation steps, message sends, and message receives at pp—ordered by the transitive closure of $\to$ restricted to $p$’s events: $e_{1}$ precedes $e_{2}$ in $H|_{p}$ whenever $e_{1}\to^{*}e_{2}$ in $H$ and both are events at $p$.
+(Causal paths through other agents’ events are not directly visible, but their ordering effects are preserved.) The frontier of $H|_{p}$ is the set of sinks of $H|_{p}$ (events with no successors under $\to$).
 Indistinguishability.
@@
 […]
-At a synchronization point, every agent sees the same admissible set (condition (iii)), and each agent knows the synchronization occurred (condition (i): $e^{*}_{j,p}$ is in its projection).
+At a synchronization point, every agent sees the same admissible set (condition (ii)), and each agent knows the synchronization occurred ($e^{*}_{j,p}$ is in its projection).
 Lemma 1 (Synchronization establishes common knowledge).
-Assume the synchronization protocol (conditions (i)–(iii)) is common knowledge among all agents. Then at each synchronization point, the current admissible set is common knowledge.
+Assume the synchronization-point conditions (i)–(ii) are common knowledge among all agents. Then at each synchronization point, the current admissible set is common knowledge.
 Proof.
-Condition (iii) gives $E(S,H)$; common knowledge of the protocol iterates this to all levels. ∎
+Condition (ii) gives $E(S,H)$; common knowledge of the conditions iterates this to all levels. ∎
 Connection to the oracle framework.
@@
 In the distributed setting, each agent has access only to a local oracle that sees its projection $H|_{p}$.
-A synchronization round (conditions (i)–(iii)) is precisely the mechanism that elevates local oracles to global ones: after synchronization, every agent’s local state includes the same admissible set (by condition (iii) and Lemma 1), so a local oracle can determine the admissible set without global access.
+A synchronization round (conditions (i)–(ii)) is precisely the mechanism that elevates local oracles to global ones: after synchronization, every agent’s local state includes the same admissible set (by condition (ii) and Lemma 1), so a local oracle can determine the admissible set without global access.
 Between synchronization points, local oracles are strictly weaker—each sees only its own projection.
@@
 In each case, the standard model silently enriches the commitment basis by assuming common knowledge of a property whose establishment would otherwise cost determination layers.
-Appendix C Proofs for Conservation and Tightness
-C.1 Conservation of sequential depth (Theorem 3)
-Let $\Phi_{0}$ be a constant-depth basis for $\mathsf{Spec}$ (Definition 11), and let $G_{\Phi_{0}}$ be its forced-dependency DAG (Definition 20).
-Each edge $\varphi\to\psi$ in $G_{\Phi_{0}}$ has a concrete witness exclusion: a triple $(H,H\cdot\varphi,o)$ such that excluding outcome $o$ at $H$ is invalid, but excluding $o$ at $H\cdot\varphi$ is valid—the dependency is witnessed by a specific outcome that cannot be excluded until $\varphi$ has taken effect.
-Any basis $\Phi$ that resolves $\mathsf{Spec}$ must also exclude each witness outcome $o$; the conservation law measures how $\Phi$ pays for these exclusions.
-Observation 4 (Semantic dependency implies circuit dependency).
-Let $\varphi_{j}$ and $\varphi_{j+1}$ be commitments from a constant-depth basis $\Phi_{0}$ such that $\varphi_{j+1}$ is forced to depend on $\varphi_{j}$.
-If a circuit $C$ computes a valid layer of a determination that achieves a witness exclusion of both $\varphi_{j}$ and $\varphi_{j+1}$, then $C$ contains a directed path from the gates computing $\varphi_{j}$’s witness exclusion to the gates computing $\varphi_{j+1}$’s witness exclusion.
-Proof.
-By forced dependency, there exist instances where $\varphi_{j+1}$’s exclusion is invalid unless $\varphi_{j}$’s exclusion has already occurred.
-If $C$ contained no path from $\varphi_{j}$’s gates to $\varphi_{j+1}$’s, then $\varphi_{j+1}$’s output would be independent of $\varphi_{j}$’s, and on such an instance the circuit would produce an invalid exclusion—a contradiction.
-∎
-Theorem 3 (Conservation law, restated).
-For any basis $\Phi$ and any valid determination over $\Phi$ that resolves $\mathsf{Spec}$ in $d$ layers with per-layer circuit depths $c_{1},\ldots,c_{d}$, $\sum_{i=1}^{d}(1+c_{i})\geq d^{*}$, where $d^{*}=\mathsf{Depth}_{\Phi_{0}}(\mathsf{Spec})$ is the determination depth under a constant-depth basis $\Phi_{0}$.
-Proof of Theorem 3.
-Let $\varphi_{1}\to\cdots\to\varphi_{p}$ be a longest path in $G_{\Phi_{0}}$, so $p=d^{*}$ (Corollary 4).
-Each edge $\varphi_{j}\to\varphi_{j+1}$ has a witness exclusion $(H_{j},H_{j}\cdot\varphi_{j},o_{j})$: outcome $o_{j}$ cannot be excluded at $H_{j}$ until $\varphi_{j}$ has taken effect.
-Fix any determination $D$ under basis $\Phi$ with $d$ layers and per-layer circuit depths $c_{1},\ldots,c_{d}$.
-Since $D$ resolves $\mathsf{Spec}$, each $o_{j}$ is eventually excluded; let $\lambda(j)$ be the $\Phi$-layer in which $o_{j}$ is first excluded.
-The richer basis $\Phi$ may exclude multiple witness outcomes within a single layer.
-For each consecutive pair $\varphi_{j}\to\varphi_{j+1}$ in the path ($j=1,\ldots,p\!-\!1$), exactly one of the following holds.
-Case 1: $\lambda(j)<\lambda(j+1)$. The predecessor’s effect is achieved in a strictly earlier $\Phi$-layer. This pair contributes at least $1$ to $d$.
-Case 2: $\lambda(j)=\lambda(j+1)=i$.
-Both effects are achieved in the same $\Phi$-layer $i$.
-By Observation 4, the circuit computing $\Phi$-layer $i$’s commitments contains a directed path from the gates for $\varphi_{j}$’s exclusion to those for $\varphi_{j+1}$’s, contributing at least $1$ to $c_{i}$.
-Case 3: $\lambda(j)>\lambda(j+1)$.
-The successor’s effect is achieved in an earlier $\Phi$-layer than the predecessor’s.
-Then $\varphi_{j+1}$’s exclusion is valid without $\varphi_{j}$’s prior exclusion, contradicting forced dependency.
-Hence Case 3 does not arise.
-Each of the p1p-1 consecutive pairs contributes at least $1$ to either $d$ (Case 1: a layer boundary) or some $c_{i}$ (Case 2: a circuit path within a layer), and each layer visited by the path contributes its $1$ term.
-Hence $\sum_{i=1}^{d}(1+c_{i})\geq p=d^{*}$.
-∎
+Appendix C Tightness Proofs
 Constrained generation instantiation.
@@
 Case 3 cannot arise for the same reason: $v_{\ell_{j+1}}$’s feasibility depends on $v_{\ell_{j}}$’s value, so the successor’s exclusion cannot be valid before the predecessor’s.
-The constrained generation task is offline: the constraint chain is fully given, so there is one complete history with no extensions.
-This rules out branch-specific conditioning—the strategy cannot make different choices for different futures, because there are no branches.
-The following remark shows why this matters.
-Remark 4 (Branch-specific commitments can bypass conservation).
-In the online setting, a basis that allows branch-specific commitments—whose effects depend on which future history materializes—can bypass the conservation bound.
-In the three-valued consensus of Example 1 (an online specification), $\mathsf{Depth}_{\Phi}(\mathsf{Spec})=2$ (under the atomic basis), but a single non-pointwise commitment can choose differently per branch (keep $b$ at $H_{1}$, $c$ at $H_{2}$, $c$ at $H_{3}$), resolving the specification in one layer with $O(1)$ circuit depth.
-The offline product-space setting of the constrained generation task rules out this escape: all commitments are applied after the history is fixed, and the coordinate basis is commutative, so no branch-specific conditioning is possible.
-C.2 Tightness of the conservation law and the three-way tradeoff
 We first prove the three-way tradeoff, then the conservation tightness.
@@
 In each case, the framework either recovers a known result with a new explanation or reveals structure that existing models do not capture.
-Determination depth arises for two reasons across these examples: enablement, where each commitment creates the sub-problem for the next layer (constrained generation, extensive-form games); and interference, where commitments that are individually valid conflict when applied in the same layer (stable matching under the rotation basis).
-Each example connects to known results in its domain: stable matching recovers Garg’s parallel algorithm [GAR20] and witnesses the orthogonality of determination and computational depth (finding a stable matching is in P, yet determination depth can be arbitrarily large); extensive-form games measure the game’s strategic depth—the moves where a player must break a tie among equally-optimal options—connecting to Selten’s trembling-hand refinement [SEL75]; chain-of-thought reasoning decomposes CoT length into determination depth, computational depth, and architectural overhead, explaining why longer chains sometimes degrade [SMA+25, SZW+25]; and BSP/LOCAL expose hidden modeling assumptions (Ameloot et al.’s non-obliviousness [ANd13]) that silently collapse determination layers.
+Determination depth arises for two reasons across these examples: commitment dependency, where each commitment creates the sub-problem for the next layer (constrained generation, extensive-form games); and forward validity, where commitments that are individually valid conflict when applied in the same layer (stable matching under the rotation basis).
+Each example connects to known results in its domain: stable matching recovers Garg’s parallel algorithm [GAR20] and witnesses the universality of determination depth (every finite depth is realized by some instance); extensive-form games measure the game’s strategic depth—the moves where a player must break a tie among equally-optimal options—connecting to Selten’s trembling-hand refinement [SEL75]; chain-of-thought reasoning decomposes CoT length into determination depth, computational depth, and architectural overhead, explaining why longer chains sometimes degrade [SMA+25, SZW+25]; and BSP/LOCAL expose hidden modeling assumptions (Ameloot et al.’s non-obliviousness [ANd13]) that silently collapse determination layers.
 The examples progress from a purely offline setting (stable matching) through online single-agent generation (chain-of-thought) to adversarial online interaction (extensive-form games) and distributed multi-agent computation (BSP and LOCAL), illustrating how determination depth provides exact diagnostics across increasing environmental complexity.
@@
 Since determination depth equals poset height (Proposition 5), stable matching realizes every possible determination depth: for every $k$, there exists an instance with determination depth exactly $k$.
-In particular, finding a stable matching is in P, yet the determination depth can be arbitrarily large—computational cost is polynomial regardless of determination depth.
 Depth-optimality of strategies.
@@
 This is precisely the parallel algorithm derived independently by Garg’s lattice-linear predicate framework [GAR20], which the determination framework recovers as a consequence of the depth characterization.
-Whether the conservation law yields new lower bounds on parallel stable matching algorithms is an open question.
 D.2 Strategic Depth in Extensive-Form Games
@@
 Proposition 6 (Strategic depth of extensive-form games).
-For a two-player extensive-form game with perfect information, the determination depth from player 1’s perspective (under the SPE-restricted specification) equals the maximum number of subgame-non-trivial player-1 nodes on any root-to-leaf path.
+For a two-player extensive-form game with perfect information, the determination depth from player 1’s perspective (under the SPE-restricted specification) is at most the maximum number of subgame-non-trivial player-1 nodes on any root-to-leaf path.
+The bound is tight whenever, between each consecutive pair of such nodes, player 2 has at least one subgame-non-trivial node.
 Proof.
@@
 Forced moves (at trivial nodes) and player-2 responses occur between commitment layers as environment events, contributing no determination depth.
-Lower bound.
+Lower bound (under the tightness condition).
 Consider two consecutive subgame-non-trivial player-1 nodes $u$ and $v$ on a root-to-leaf path, with $v$ deeper than $u$.
@@
 The distinction is analogous to the BSP diagnostic (Section D.4): just as BSP overcharges by counting communication rounds that resolve no ambiguity, game-tree depth overcharges by counting moves that present no relational choice.
-The analysis is restricted to finite perfect-information games (Zermelo’s setting); extending to imperfect information requires handling information sets and is left open.
+The analysis is restricted to finite perfect-information games (Zermelo’s setting [ZER13]); extending to imperfect information requires handling information sets and is left open.
 Remark 5 (Error amplification under bounded rationality).
@@
 An indirect mechanism (e.g., an ascending auction [MIL00]) trades strategic depth for communication: more rounds of bidding, but less information per round.
-This echoes the depth–width–communication tradeoff (Theorem 2): formalizing the precise relationship—whether indirect mechanisms obey an analogous conservation law trading determination depth against per-round communication and outcome-set width—is an open problem.
+Whether indirect mechanisms obey a conservation law analogous to Theorem 3—trading determination depth against per-round communication and outcome-set width—is an open problem.
 D.3 Decomposing Chain-of-Thought Length
@@
 The constrained generation task (Section 3) serves as the witness because it isolates determination depth from computational depth: each commitment is a constant-time table lookup, so the sequential cost is entirely semantic.
-The total CoT length $T$ is lower-bounded by three formally independent quantities: (i) determination depth dd—non-commuting commitment layers, witnessed by the constrained generation task (Section 3); (ii) computational depth cc—inherent per-step sequentiality from bounded transformer depth (TC0-like [FZG+23]), witnessed by graph connectivity [MAA+25]; and (iii) architectural overhead—tokens spent on information management (context summarization, backtracking, re-derivation), reducible by changing the architecture without changing the task.
+The total CoT length $T$ is lower-bounded by three independent quantities: (i) determination depth dd—non-commuting commitment layers, witnessed by the constrained generation task (Section 3); (ii) computational depth cc—inherent per-step sequentiality from bounded transformer depth (TC0-like [FZG+23]), witnessed by graph connectivity [MAA+25]; and (iii) architectural overhead—tokens spent on information management (context summarization, backtracking, re-derivation), reducible by changing the architecture without changing the task.
 Proposition 7 (CoT lower bound from determination depth).
@@
 Appendix F Related Work
-This section positions determination depth among existing notions of staged computation and semantic dependency.
-Closest antecedents treat staging either as an operational resource (rounds, adaptivity, oracle calls) or as a syntactic discipline (stratification, fixpoint nesting).
-Our contribution is to identify staging as a semantic invariant of resolution—defined prior to any particular model or language.
+This section positions determination depth relative to existing complexity measures that involve sequential staging.
+These measures treat staging either as an operational resource (rounds, adaptivity, oracle calls) or as a syntactic discipline (stratification, fixpoint nesting); determination depth differs in that it measures the semantic cost of irrevocable commitment, independent of any particular model or language.
 Parallelism, depth, and inherent sequentiality.
@@
 Determination depth measures the inherent sequentiality of choosing an outcome from a relation: commitment dependencies force some decisions to wait for others.
-The two are orthogonal: a function has determination depth 0 regardless of its circuit depth, while a relation can have high determination depth even when every individual commitment is a constant-depth circuit (Section 3).
-This orthogonality is witnessed in the offline setting by the constrained generation task (Section 3); in the online setting, forward validity constraints add a further source of determination depth that circuit models do not capture at all.
+The two are distinct: a function has determination depth 0 regardless of its circuit depth, while a relation can have high determination depth even when every individual commitment is a constant-depth circuit (Section 3).
+In the online setting, forward validity constraints add a further source of determination depth that circuit models do not capture at all.
 PRAM and BSP models similarly separate local computation from global synchronization, charging depth to barriers or rounds [BC82, VAL90].
 In the online (distributed) setting, these models can both undercharge (when a round boundary hides an irrevocable commitment that the model treats as primitive) and overcharge (when communication structure is conflated with semantic resolution); Section D.4 develops this diagnostic in detail.
-More broadly, classical hierarchy theorems establish strict separations based on time, space, or alternation depth [SIP83]; our determination depth hierarchies are orthogonal, arising from semantic dependency rather than resource bounds.
+More broadly, classical hierarchy theorems establish strict separations based on time, space, or alternation depth [SIP83]; our hierarchies arise from semantic dependency rather than resource bounds.
 Trace monoids and partial commutation.
@@
 In determination theory, commutation is dynamic: whether two commitments commute depends on the current refined specification, and applying one commitment can create or destroy independence among the remaining ones.
-This dynamic commutation arises for different reasons in the two settings: in the online setting, forward validity constraints (a commitment that is safe alone may become unsafe after another commitment narrows the admissible set at some extension) create and destroy independence; in the offline setting, enablement (committing to one position determines which choices remain feasible at the next) produces the same effect.
+This dynamic commutation arises for different reasons in the two settings: in the online setting, forward validity constraints (a commitment that is safe alone may become unsafe after another commitment narrows the admissible set at some extension) create and destroy independence; in the offline setting, commitment dependency (committing to one position determines which choices remain feasible at the next) produces the same effect.
 Second, this state-dependence makes the analogue of the Foata normal form non-unique—different layering choices lead to different refined specifications and potentially different minimum determination depths—and computing the minimum height becomes NP-hard (Proposition 8), in contrast to the polynomial-time computability of static Foata height.
@@
 Adaptivity depth measures how many times a strategy must look; determination depth measures how many times it must choose.
-The two are orthogonal (Observation 3, Section 3): pointer chasing at sparsity $s=1$ has high adaptivity depth but determination depth zero (the answer is unique); constrained generation in the offline setting has determination depth $k$ but adaptivity depth zero (the entire input is given).
+The two are independent (Observation 3, Section 3): pointer chasing at sparsity $s=1$ has high adaptivity depth but determination depth zero (the answer is unique); constrained generation in the offline setting has determination depth $k$ but adaptivity depth zero (the entire input is given).
 Round hierarchies in communication complexity.
@@
 Layered semantics appear in database theory through stratified negation, well-founded semantics, and alternating fixpoints [GRS91]; in descriptive complexity, fixpoint nesting and alternation depth classify expressive power [IMM99].
-These layerings are properties of a program under a chosen semantics: a stratified Datalog program with $k$ strata requires $k$ non-commuting sealing commitments (one per stratum), but applying well-founded or stable semantics to the same program yields a different commitment structure and potentially a different determination depth.
-Determination depth, by contrast, is a property of the specification (the mapping from histories to admissible outcome sets), independent of any particular logical formalism or semantics used to define it.
-Two logic programs that define the same specification have the same determination depth, even if they differ in strata count, fixpoint nesting, or use of negation.
-Conversely, applying different negation semantics to the same program can yield different specifications and hence different determination depths—a distinction that strata counts alone cannot express.
+These layerings are properties of a program under a chosen semantics.
+Determination depth, by contrast, is a property of the specification (the mapping from histories to admissible outcome sets), independent of any particular logical formalism or semantics: two programs that define the same specification have the same determination depth, even if they differ in strata count, fixpoint nesting, or use of negation.
+The distinction matters because different negation semantics applied to the same program can yield different specifications and hence different determination depths.
+Each semantics induces a different commitment basis—stratified semantics requires a linear chain of sealing commitments, well-founded semantics an alternating sequence, stable semantics a shallow branching among incompatible determinations—a distinction strata counts alone cannot express.
 Leaf languages.
@@
 More broadly, the leaf language program provided an elegant classification of existing complexity classes but did not yield new separations or lower bounds.
-Our framework is aimed at a different target: not reclassifying known classes, but identifying an axis of complexity orthogonal to computation (Section 3), producing new tradeoffs (Theorems 12), and diagnosing existing models (Sections $D.4$).
+Our framework is aimed at a different target: not reclassifying known classes, but identifying an axis of complexity distinct from computation (Section 3), producing new tradeoffs (Theorems 12), and diagnosing existing models (Sections $D.4$).
 Coordination and monotonicity.
@@
 These results concern computational depth—the inherent sequentiality of evaluating a function.
-Our exponential separation (Section 3) establishes an orthogonal source of sequential advantage: determination depth, arising from non-commuting commitments in relational specifications.
+Our exponential separation (Section 3) establishes a distinct source of sequential advantage: determination depth, arising from non-commuting commitments in relational specifications.
 The separation is an offline result (the full constraint chain is given); the decomposition of Appendix D.3 extends to the online autoregressive setting, unifying computational depth, determination depth, and architectural overhead into a single framework.
@@
 Across these domains, notions of depth have appeared as proxies for sequentiality, staging, or adaptivity.
-Our contribution is to identify determination depth and determination cost as semantic complexity measures—arising from forward validity constraints in the online setting and from commitment dependency in the offline setting—that are orthogonal to computational complexity.
+Our contribution is to identify determination depth and determination cost as semantic complexity measures—arising from forward validity constraints in the online setting and from commitment dependency in the offline setting—that are distinct from computational complexity.
 Determination depth measures the minimum sequential layers of irrevocable commitment; determination cost measures the total number of commitments; the exponential separation (Theorem 1) shows that reducing one without increasing the other is impossible.
@@
 We used Claude (Anthropic), ChatGPT (OpenAI), and Gemini (Google) to assist with exposition, structural organization, proof review, and adversarial feedback throughout the manuscript.
-The tools materially affected the prose and presentation in all sections, including section ordering, framing of arguments, and selection of open problems.
+The tools materially affected the prose and presentation in all sections, including suggestions for tone and framing of arguments, and potential applications and open problems.
 All mathematical content, definitions, theorems, and proofs are the authors’ own.